### v1

# Layer-2 Arbitrage: An Empirical Analysis of Swap Dynamics and Price Disparities on Rollups

Abstract
This paper explores the dynamics of Decentralized Finance (DeFi) within the Layer-2 ecosystem, focusing on Automated Market Makers (AMM) and arbitrage on Ethereum rollups. We observe significant shifts in trading activity from Ethereum to rollups, with swaps on rollups happening 2–3 times more often, though, with lower trade volume. By examining the price differences between AMMs and centralized exchanges, we discover over 0.5 million unexploited arbitrage opportunities on rollups. Remarkably, we observe that these opportunities last, on average, 10 to 20 blocks, requiring adjustments to the LVR metrics to avoid double-counting arbitrage. Our results show that arbitrage in Arbitrum, Base, and Optimism pools ranges from 0.03% to 0.05% of trading volume, while in zkSync Era it oscillates around 0.25%, with the LVR metric overestimating arbitrage by a factor of five. Rollups offer not only lower gas fees, but also provide faster block production, leading to significant differences compared to the trading and arbitrage dynamics of Ethereum.
Keywords: Arbitrage, Automated Market Maker, Rollup, MEV
1 Introduction
Ethereum has undergone a series of substantial enhancements aimed at increasing its functionality, addressing scalability challenges, and improving user interactions [2]. A pivotal upgrade was the transition from Proof of Work (PoW) to Proof of Stake (PoS), referred to as The Merge, which occurred throughout 2022 and 2023. After that, the most significant milestone was the Dencun upgrade, implemented on March 13th, 2024. It introduced Proto-Danksharding, specifically designed for rollups: Layer-2 (L2) scaling solutions. This upgrade marked the initial phase of Ethereum’s rollup-centric roadmap and introduced blobs, temporary storage for rollup transaction data, streamlining verification, and reducing data storage costs [2]. Consequently, the gas fees on rollups have been significantly reduced, as shown in Figure 1.
Figure 1: Median transaction fees, measured in Gwei, for selected Ethereum rollups, pre- and post-Dencun Upgrade on March 13th, 2024.
The reduction in gas fees has influenced the use of Decentralized Finance (DeFi) on rollups. In recent years, the Total Value Locked (TVL) and the trading volume on rollups have exhibited a notable increase, with certain rollups now surpassing Ethereum in daily transaction numbers [1, 3]. Furthermore, the dynamics of L2 DeFi is not solely altered by reduced gas fees. The second important factor is the faster block production, typically less than 2 seconds in contrast to Ethereum’s 12 seconds [28]. The combination of lower gas fees and faster block production enables arbitrageurs to capitalize on smaller price discrepancies between trading venues more quickly. Although arbitrage transactions are performed predominantly on Ethereum through Maximal Extractable Value (MEV) Boost Auctions [16], an equivalent MEV boost mechanism is not present in rollups. The absence of MEV boost auctions on rollups can be explained by the fact that most rollups rely on centralized sequencers to aggregate transactions into L1 batches.
The centralized sequencers and the associated censorship risk remain a concern within the L2 DeFi ecosystem. Additional concerns encompass transaction finality, the time period until an L2 transaction becomes irrevocable from the rollup (soft finality) or the underlying layer-1 blockchain (hard finality) [31]. Rollups can be categorized into optimistic and Zero-Knowledge (ZK) rollups based on transaction validity and proof mechanisms [11]. Despite the seven-day hard finality period, optimistic rollups have gained user adoption more rapidly due to better Ethereum Virtual Machine (EVM) compatibility. However, ZK-rollups offer higher transaction compression, throughput, and faster hard finality, ranging from minutes to hours. The aforementioned factors — reduced gas fees, absence of MEV auctions, varying finality speeds, and accelerated block production — contribute to distinct DeFi dynamics on rollups, which warrant further investigation.
With the emergence of decentralized rollup sequencers [26], there arises a potential risk that adverse practices associated with MEV on Ethereum [27] transfer to the L2 DeFi ecosystem. To address this concern, a time-boost mechanism within sequencers has been proposed [22]. This mechanism is specifically designed to allow back-running MEVs while preventing front-running activities. Front-running MEV can facilitate various forms of adversarial actions, such as sandwich attacks on traders [19] and Just-In-Time (JIT) liquidity attacks on Liquidity Providers (LPs) within AMM-DEXs [32]. In contrast, back-running MEV enables immediate arbitrage at DEXs or liquidations at lending protocols. Faster arbitrage execution benefits traders, as it eliminates price discrepancies between trading venues. Additionally, MEV time-boost auctions can generate new revenue for sequencer operators.
This paper contributes to the research on L2 DeFi by empirically examining the differential swap dynamics on rollups and Ethereum, considering the factors mentioned above, such as lower gas fees, faster block production, and the absence of MEV auctions. Furthermore, we quantify the potential arbitrage opportunities across rollups and between rollups and CEXs that can be the basis for the revenue estimation from the time-boost MEV auctions in the rollup sequencers.
Related Work.
There exists a substantial body of research on DeFi, particularly focused on AMM-DEXs. In particular, the systematization of knowledge (SoK) by Xu et al. [30] and empirical investigations of Uniswap (v3) on the Ethereum blockchain [18, 9, 17, 21, 12] are prominent. Most of these studies evaluate the profitability of LPs, specifically addressing impermanent loss. The analysis by Adams et al. [4] delves into the costs and dynamics of swaps on Uniswap (v3) on Ethereum. Our research extends this by breaking down swap fees on rollups and assessing the influence of blobs and block time on swap dynamics. Our findings are consistent with the simulated decomposition of swap fees in rollups presented in Gogol et al. [14].
The loss versus rebalancing metric (LVR), which compares arbitrageur profits with those of LPs, was introduced by Milionis et al. [24, 25] and empirically measured by Fritsch et al. [10]. LVR evaluates the LP portfolio against a rebalancing portfolio that perfectly hedges the market risk of the LP position. Our investigation quantifies the absolute profit of arbitrageurs, also known as the net LVR (per block), incorporating significant adjustments. First, LVR assumes that the price discrepancies between AMMs and target exchanges, such as CEXs, disappear after each block. Our findings confirm this assumption for Uniswap (v3) on Ethereum; nonetheless, the decay period of price discrepancies on rollups extends over multiple blocks. Our methodology adjusts the LVR to this scenario and prevents double counting of the same arbitrage opportunities. Second, LVR and the rebalancing portfolio assess the LP’s loss relative to arbitrageurs, thereby containing impermanent loss. Our metrics measure the absolute profit of the arbitrageur. Practically, arbitrageurs do not replicate the rebalancing portfolio but instead capitalize on arbitrage opportunities [16].
The persistence of price disparities on AMMs within zkSync Era, a ZK rollup, was initially reported by Gogol et al. [13]. Our paper builds on that research by providing: i) a comprehensive analysis of rollups (Arbitrum, Base, Optimism, in addition to zkSync Era) alongside Ethereum, ii) enhanced precision in market price accuracy and per-block analysis, (improving from a 1-minute to a 1-second resolution), and iii) an in-depth examination of Uniswap (v3), focusing on the most actively traded pools.
Heimbach et al. [16] focus on the estimation of arbitrage executed through MEV auctions on Ethereum, specifically addressing non-atomic arbitrage opportunities between AMMs and CEXs. Our research extends this investigation by examining rollups, where MEV auctions have not yet been implemented. Similarly, we establish the minimum threshold for arbitrage opportunities and derive the arbitrage value using AMM formulas. However, this work seeks to quantify the unexploited arbitrage, which occasionally persisted for several minutes, in contrast to the arbitrage realized via MEV on Ethereum. Other studies on arbitrage [7, 20, 8] focus primarily on AMM-DEX on Ethereum and/or cyclic arbitrage opportunities.
Torres et al. [29] is the first effort to quantify MEV in roll-ups. Despite the absence of MEV auctions, this research examines the heuristic for back-running arbitrage MEV, specifically focusing on swap events occurring within the same transaction in the block, within the context of cylindrical arbitrage. Our objective is to assess the potential for non-atomic arbitrage, involving CEXs, or cross-rollup price discrepancies.
Mamageishvili et al.  [22] introduce a time boost mechanism within rollup sequencers. This mechanism facilitates back-running MEV while preventing front-running MEV. In subsequent work [23], a shared sequencer was proposed, enabling cross-rollup MEV auctions, such as those for arbitrage exploitation. However, these studies lack empirical quantification of arbitrage opportunities within rollups, a gap that this work seeks to address.
Contributions.
This paper advances DeFi research within rollups. We conducted an empirical investigation into the distinct swap dynamics in optimistic and ZK rollups, as compared to Ethereum, taking into account specific rollup characteristics, such as lower gas fees, faster block production, and the absence of MEV auctions. Subsequently, we analyze the arbitrage opportunities on rollups that can be exploited by arbitrageurs, thereby providing new revenue streams for sequencers offering back-running MEV auctions. Our principal contributions are as follows.
  • Examination of Arbitrage and Swapping Costs on Rollups: We conduct a comparative study of the costs of swapping and capturing arbitrage on rollups compared to Ethereum. Our findings indicate that, unlike Ethereum, transactions on rollups occur 2–3 times more frequently but are smaller in size. Furthermore, swaps on rollups are generally not impacted by front- or back-running block slippage typically caused by MEV. Block times impact rollup transactions; Ethereum averages one swap per block, while rollups have fewer swaps per block despite higher total swaps. Specifically, swaps occur every third block on Base, fifth on Optimism, and tenth on Arbitrum.
  • Decay of Price Disparities: We show that price differences last for 10–20 blocks in rollups, necessitating modifications in the empirical assessment of LVR metrics. We present a framework for empirically estimating arbitrage in cases where the decay period of price differences spans multiple blocks, capturing only the maximum arbitrage value (MAV) during prolonged price disparities. The empirical MAV method, which filters out repeated arbitrage, indicates that LVR metrics have overestimated arbitrage by a factor of five.
  • Estimation of Arbitrage on Rollups: By examining the price differences between Uniswap (v3) and Binance, we discovered over 0.5 million unexploited arbitrage opportunities on rollups. The MAV metric on Arbitrum, Base, and Optimism pools varies from 0.03% to 0.05% of trading volume, whereas in zkSync Era, it is approximately 0.25%, highlighting significant unresolved price differences.
Paper Organization.
Section 2 offers foundational knowledge on rollups and AMMs. Section 3 details the dataset utilized in our analysis, while Section 4 breaks down swap costs on Ethereum rollups, both pre- and post-Decun (blobs). Section 5 sets the theoretical foundation for arbitrage in AMMs and its estimation framework in rollups. Section 6 presents the results of our empirical investigations. Section 7 examines the significance of arbitrage in rollups and proposes future research directions. Lastly, Section 8 summarizes our findings as well as some final remarks.
2 Background
This section introduces rollups, a new mechanism of layer-2 blockchain scaling, and formalizes the blockchain terminology used throughout the manuscript. Furthermore, it presents the basics of Decentralized Exchanges (DEXs) and Automated Market Makers (AMMs).
2.1 Layer-2 Blockchains
There are two main approaches — Layer-1 (L1) and Layer-2 (L2) scaling — to tackle blockchain scalability challenges and the associated increases in gas fees. L1 scaling involves creating a new blockchain with its own consensus mechanisms and a unique physical infrastructure to support the network. In contrast, L2 scaling uses a different method: it carries out computations off the main blockchain while recording their results on the underlying chain. The main L2 scaling solutions include state (payment) channels, plasma, validium, and rollups. While plasma and state channels aim to move both data and computation off-chain, rollups and validium are non-custodial: they shift computation off-chain but keep compressed transaction data (rollups) or validity proofs (validiums) on the underlying L1 chain. As a result, rollups and validium require fewer trust assumptions compared to state channels, plasma chains, or L1 networks, with data security and integrity being ensured by the security of the underlying chain.
Rollup.
A rollup functions as a virtual blockchain, producing transaction blocks and recording them on the main chain as compressed batches. However, it operates with the risk of adversarial actions such as stopping block production, generating invalid blocks, withholding data, or engaging in other malicious activities. To ensure the correctness of the transaction state in these untrusted environments, rollups use one of two approaches: the optimistic approach and zero-knowledge (ZK) proofs. Figure 2 illustrates the rollup architecture, highlighting its principal components: sequencers and verifiers. Sequencers consolidate transactions to be rolled up onto the L1 chain, thereby achieving gas fee reductions. Verifiers, which are smart contracts on L1, validate the transaction block created by the sequencer, thus ensuring its correctness. Furthermore, ZK-rollups utilize provers responsible for generating ZK validity proofs.
Figure 2: Architecture of a rollup
Soft and Hard Finality.
Transaction finality refers to the point at which the state of a transaction in the blockchain ledger becomes permanent. Within the context of rollups, there are two types of transaction finality: hard finality, which is the permanent state of a transaction on the L1 chain, and soft finality, which is the irreversible state of a transaction on the L2 chain. [31]
Dencun Upgrade and Blobs.
On 13 March 2024, Ethereum underwent the Dencun Upgrade, which introduced temporary data storage known as “blobs” [2]. Blobs are designed to enhance the efficiency of Ethereum rollups. Ethereum nodes store blob data for 18 days, during which rollups validate their transactions. After this 18-day period, the blobs are removed from the Ethereum blockchain. Rollups are not required to keep transaction details from blobs; however, they typically store off-chain all transaction data for a better user experience.
2.2 Automated Market Makers
Decentralized Exchanges (DEXs) facilitate the direct exchange of tokens on the blockchain in a non-custodial and atomic way. The implementation of a Limit Order Book (LOD) on a blockchain is impractical due to storage costs and computational complexity. Consequently, DEXs employ Automated Market Makers (AMMs) to determine the exchange rate between tokens. This exchange rate is derived based on the quantity of tokens held in liquidity pools, which are supplied by Liquidity Providers (LPs).
Constant Function Market Makers (CFMMs) are the most common type of AMMs. They use a reserve curve function $F:\mathbb{R}^{N}_{+}\rightarrow\mathbb{R}$, which maps $N$ token reserves $x_{i}$ to a fixed invariant $L$. The reserve curve functions are carefully designed to improve capital efficiency within liquidity pools and to reduce the price impact for traders [30]. Each token trade on an AMM involves a trading fee, usually a fixed percentage of the swap amount, paid by the trader. These fees, or a portion of them, are then distributed proportionally among the LPs in the pool as compensation for their contribution. For simplicity, we refer to them as LP fees.
Constant Product Market Makers (CPMM).
Uniswap (v2), one of the first and most popular AMMs, operates as a CPMM [5]. Its trading mechanism is governed by the formula:
$x_{1}\cdot x_{2}=L,$
where $x_{i}$ represent the balances of the two tokens in the liquidity pool, and $L$ is a constant indicating the pool’s liquidity level.
CPMM with Concentrated Liquidity (CLMM).
Uniswap v3 introduced concentrated liquidity, an enhancement to CPMMs, to improve the efficiency of the utilization of liquidity within the pool [6]. This allows LPs to specify a price range to provide liquidity. For a two-token pool ($N=2$) within the price range $[p_{j},p_{k}]$, the trade invariant is:
$\left(x_{1}+\frac{L}{\sqrt{p_{k}}}\right)\left(x_{2}+L\cdot\sqrt{p_{j}}\right)=L^{2},$
where the variables hold similar meanings as previously described. The downside of CLMMs is that LPs must continuously monitor and adjust the price range of their liquidity provision, as only trades within this range earn LP fees. Thus, LPs require liquidity provision strategies to manage LP positions and effectively minimize gas fees.
Spot Price.
The spot exchange rate between tokens can be determined as the slope of the $x_{1}-x_{2}$ curve, illustrated in equations 2.2 and 2.2), with partial derivatives of the curve function $F$.
2.3 Cost of Swapping
The total cost of executing swap transactions on an AMM includes gas costs and explicit and implicit DEX fees [30, 15]. For clarity, we express this as follows.
[…]
Gas Fee.
Unlike gas fees on L1 blockchains, the gas fees for rollup transactions are not predetermined as they consist of two parts: those charged by the sequencer and those charged by the underlying L1 network. The sequencer estimates the gas fee for a transaction, and this estimated amount is deducted from the originating address. Later, once the transaction, along with others, is posted on the L1 chain, the exact gas fees from L1 are determined. Depending on the specific rollup implementations, any overpayment by the transaction originator is either refunded immediately (e.g., zkSync Era) or during the next transaction (e.g., Arbitrum).
LP Fee.
Also known as trading fees, LP fees are the explicit fees paid by the trader to the DEX to swap tokens. They typically range between 1 and 50 basis points (0.01% and 0.05%) of the swap volume, and are fully or partially redistributed to LPs.
Block Slippage.
This represents an implicit swap cost when trading on an AMM and can result in unexpected changes in the execution price due to the order of transactions within a block. The impact of slippage can be either positive or negative, depending on whether it benefits or disadvantages the trader in terms of the executed price.
Price Impact.
Price impact is the effect of an individual trade on the market price of the underlying asset pair. It is directly related to the size of the trade and the amount of liquidity available in the AMM’s pool.
3 Data Collection
We analyze swap data from Ethereum and its roll-ups with the highest trading volumes. This includes Arbitrum, Base, and Optimism (optimistic rollups), as well as zkSync Era (ZK-rollup). For each blockchain, we focus on the most actively traded pool, WETH-USDC on Uniswap (v3), during the period from December 31st, 2023, to April 30th, 2024. Although the Uniswap (v3) deployment on zkSync Era does not have a graphical user interface, users can interact with it via DEX aggregators such as Oku Analytics.
The data set is sourced from blockchain archive nodes provided by Nansen. The number of analyzed swaps, transactions, and block ranges is detailed in Table 1. Using the event logs of the Swap method, we recalculate the historical spot prices and liquidity in ticks within the USDC-WETH pools on Uniswap (v3). For Uniswap (v3), the spot price after the swap is obtained from sqrtPriceX96 in the event logs. Market data for the ETH-USDC exchange rate on CEX are obtained from Binance. We examine 1-second Binance closing prices, noting that block production on Ethereum occurs every 12 seconds, while it is around 1–2 seconds on Base, Optimism, and zkSync Era. On Arbitrum, block production can be as fast as 0.25 seconds. Thus, for each blockchain, except for Arbitrum, computations are performed at the block level. For Arbitrum, we analyze the last block in a second. The spot price on DEX is determined by the last swap within the block.
Chain Swaps Transactions Blocks Block Range Ethereum $761005$ $749818$ $475409$ 1890889618908896189088961890889619771559197715591977155919771559 Arbitrum $2400000$ $2367361$ $1709619$ 187373628187373628187373628187373628206540031206540031206540031206540031 Arbitrum (USDC.e) $2258469$ $2232543$ $1648839$ 165788868165788868165788868165788868206540037206540037206540037206540037 Base $1687530$ $1636196$ $1145789$ 863892986389298638929863892913866123138661231386612313866123 Optimism (USDC.e) $1186780$ $1136839$ $902789$ 114234215114234215114234215114234215119461410119461410119461410119461410 zkSync Era (USDC.e) $46417$ $46379$ $45364$ 2290992322909923229099232290992332843035328430353284303532843035
Table 1: List of analyzed WETH-USDC and WETH-USDC.e liquidity pools on Uniswap (v3) on Ethereum and its rollups, spanning the period from 31st December 2023 to 30th April 2024.
Native USDC and Bridged USDC.e.
USDC is a fiat-backed stablecoin issued by Circle. It can exist on rollups either as a native token or a bridged one. The native USDC is directly supported by Circle’s off-chain reserves, while USDC.e is a token bridged from Ethereum. Generally, USDC.e is introduced to the blockchain first, followed by the issuance of the native token. In most rollups, USDC and USDC.e coexist for legacy purposes. The native USDC is resistant to hacking and can be easily transferred across blockchains using the Circle bridge.
4 Swap Dynamics
Tables 2 and 3 present the characteristics of swaps on Ethereum and its rollups. As shown in Table 2, swap transactions on rollups have a lower volume compared to those on Ethereum, although they occur more frequently. On average, there are 2–3 times more swaps on rollups, but the trading volume is about five times lower. Consequently, the median swap volume on Ethereum is $4239.45$ USD, while it is $2201.23$ USD on Arbitrum, $173.46$ USD on Base, $331.28$ USD on Optimism, and only $20$ USD on zkSync Era.
Table 3 explains the effect of block times on rollup swap transactions. Despite having a higher number of swaps, there are fewer swaps per block on rollups because of faster block production. In Ethereum, there are, on average, $0.88$ swaps in every block, with every block containing swap transactions having 1.6 swap events. In contrast, every third block on Base, fifth block on Optimism, and tenth block on Arbitrum contain a swap. Furthermore, there are fewer swap events per block with swap transactions (approximately 1–1.47) on rollups.
On average, there are slightly more swap events per transaction (1.01) on both rollups and Ethereum, indicating that some swaps are performed in batches. This could imply the exploitation of MEV by certain wallets or decentralized exchange (DEX) aggregators initiating swaps on behalf of users.
Chain Swaps Total Volume mn Avg Swap Volume Std Swap Volume Median Swap Volume
Ethereum $761005$ $35175$ $46222.26$ $142890.62$ $4239.45$
Arbitrum $2400000$ $7266$ $3027.66$ $6738.06$ $1125.01$
Arbitrum (USDC.e) $2258469$ $8159$ $3612.86$ $5976.01$ $2201.23$
Base $1687530$ $2996$ $1775.55$ $5611.25$ $173.46$
Optimism (USDc.e) $1186780$ $1457$ $1228.30$ $2842.73$ $331.28$
zkSync (USDC.e) $46417$ $8.01$ $172.67$ $426.52$ $20.99$
Table 2: An analysis of the absolute swap volumes within the WETH-USDC and WETH-USDC.e liquidity pools on Uniswap (v3) across the Ethereum network and its rollups, spanning the period from 31st December 2023 to 30th April 2024.
Chain Block Time (s) Swaps per Block Swaps per Transaction Swaps Per Block with Transaction Ethereum 12.12 $0.88$ $1.01$ $1.60$ Arbitrum 0.25 $0.13$ $1.01$ $1.40$ Arbitrum (USDC.e) 0.25 $0.06$ $1.01$ $1.37$ Base 2.00 $0.32$ $1.03$ $1.47$ Optimism (USDC.e) 2.00 $0.23$ $1.04$ $1.31$ zkSync Era (USDC.e) 1.05 $0.00$ $1.00$ $1.02$
Table 3: An analysis of the swaps per transactions and blocks within the WETH-USDC and WETH-USDC.e liquidity pools on Uniswap (v3) on the Ethereum network and its rollups, spanning the period from 31st December 2023 to 30th April 2024.
5 Theory of Arbitrage and LVR
Price disparities across different trading platforms (both DEXs and CEXs) present arbitrage opportunities. In this section, we calculate the transaction volume required to align prices between AMM-DEX and CEX, given the existing token reserves in an AMM liquidity pool. Trading this volume maximizes the arbitrageur’s profit, which is called net LVR [24], or the Maximal Arbitrage Value (MAV) [13]. For simplicity, we further assume that trading on the CEX has no price impact due to the typically higher reserves compared to AMM-DEXs.
Maximal Arbitrage Value.
Consider a scenario in which there is a price discrepancy between a CEX and an AMM, the CEX having a substantially higher market depth, resulting in zero market impact. Let $P_{c}$ and $P_{a}$ be the prices at CEX and AMM, respectively, and assume that $P_{c}$ < $P_{a}$. Let $V_{max}$ denote the maximum trade volume that was bought on the CEX for $P_{c}$ and sold on the AMM for $P_{a}$ equals the prices. With $V_{max}$ being the trading volume that equals the prices at AMM and CEX and percentage price impact $\rho(V_{max})$, MAV is defined as:
$MAV=V_{max}\cdot\Big{(}P_{a}\cdot(1-\rho(V_{max}))-P_{c}\Big{)},$
or equivalently
$MAV=V_{max}\cdot(P_{a}-P_{c})-V_{max}\cdot P_{a}\cdot\rho(V_{max}).$ (1)
Arbitrage between a CPMM and CEX.
Let us assume that we have two tokens $X$ and $Y$, the respective reserves in the liquidity pool of AMM are $x_{1}=x$ and $x_{2}=y$. Assuming that we are dealing with a CPMM, the act of trading follows the $x_{1}\cdot x_{2}=x\cdot y=L$ formula, and the instantaneous exchange rate $P_{a}$ for an infinitesimal trade is
$P_{a}=\frac{x}{y}.$ (2)
and the percentage price impact on a CPMM (such as Uniswap v2) is given by $\Delta x/x$ for a swap of $\Delta x$ for $\Delta y$. the MAV can be expressed as:
$MAV=V_{max}\cdot(P_{a}-P_{c})-\frac{V_{max}^{2}P_{a}}{x},$ (3)
Following [16, 13] we can compute the first derivative of Eq. (3) with respect to $V_{max}$ to find the max points, receiving:
$V_{max}=\frac{y\cdot(P_{a}-P_{c})}{2P_{a}}.$ (4)
And, by substituting (4) into (3) the MAV is written as:
$MAV=\frac{y\cdot(P_{a}-P_{c})^{2}}{4P_{a}}.$ (5)
Arbitrage between a CLMM and CEX.
Following [13], we generalize the MAV formula of CPMM - Uniswap (v3) for concentrated liquidy in Uniswap (v3). The conservation function of a Uniswap v3 pool aggregates individual LPs’ conservation functions for different ticks, each dependent on the exchange rate range specified by the LP.
Ticks occur at each price $p(i)=1.0001^{i},i\in[1,2,...]$. Suppose that an LP offers liquidity ($x,y$) exclusively to users trading within a specific range of exchange rates defined by two surrounding consecutive ticks, i.e. $[P_{a}/\alpha,P_{a}\alpha]$ with $\alpha>1$ and $P_{a}=\frac{x}{y}$ as the current spot price. By definition, there exists some $i$ such that $[P_{a}/\alpha,P_{a}\alpha]\rightarrow[p(i),p(i+1)]$, and thus we denote for the sake of clarity ($x,y$) as ($x_{i},y_{i}$). Within this setting, the shape of the trading function is identical to the case of liquidity provision of equivalent reserves
$x_{i}^{equiv}=\frac{x_{i}}{1-1/\sqrt{\alpha}},\text{ and }y_{i}^{equiv}=\frac{y_{i}}{1-1/\sqrt{\alpha}}$ (6)
under Uniswap v2.
We now allow the proportion of reserves to vary within the price range defined by the two ticks under consideration, and for simplicity, define $r_{i,x},r_{i,y}\geq 0$ as such varying quantities. Clearly, there is an upper bound to the change in reserves due to the implied movement to the next adjacent tick bounds. This translates into $0\leq r_{i,x}\leq x_{i}\cdot(\sqrt{\alpha}+1)$ and $0\leq r_{i,y}\leq y_{i}\cdot(\sqrt{\alpha}+1)^{2}$. Thus, we can derive (similarly as for Uniswap v2) the percentage price impact when trading within two ticks $[p(i),p(i+1)]$ on Uniswap v3 [30]. This is
$\rho_{i}(\Delta y)=\frac{\Delta x}{r_{i,x}+\frac{x_{i}}{\sqrt{\alpha}-1}},$ (7)
with the constraint that $\Delta y\leq r_{i,y}$, since otherwise we deplete one of the two reserves and move to the next tick. Thus, our MAV formula becomes
$MAV_{i}=V_{i,max}\cdot(P_{a}-P_{c})-V_{max}\cdot P_{a}\cdot\rho_{i}(V_{i,max}),$ (8)
where $p(i)\leq P_{c}\leq p(i+1)$, and which needs to be solved enforcing the constraint $\Delta y\leq r_{i,y}$.
If the price misalignment spans multiple ticks, one then needs to iteratively compute $MAV_{i}$ and sum the related profits until realignment, i.e.
$MAV_{Uniswap\_v3}=\sum_{i:P_{a}\in[p(i),p(i+1)]}^{i:P_{c}\in[p(i),p(i+1)]}MAV_{i}.$ (9)
6 Empirical Results
When price discrepancies arise between trading platforms, arbitrageurs can take advantage of these opportunities to profit from price corrections. This activity benefits both the arbitrageurs, who gain from the price differences, and traders, who enjoy promptly adjusted and accurate exchange rates. The only parties that might suffer losses from price corrections are LPs at AMM-DEX. As a result, arbitrage profits are compared with LP losses in the loss-versus-rebalancing (LVR) metric [24].
In this section, we examine the price differences between WETH-USDC (and USDC.e) AMM pools on Ethereum and its rollups — Arbitrum, Base, Optimism, zkSync Era — and Binance, the leading cryptocurrency CEX. We measure the decay time of these discrepancies and quantify the resulting arbitrage opportunities using the proposed Maximal Arbitrage Value (MAV) metric, similar to the net LVR. We begin by presenting a methodology to estimate the empirical MAV, which accounts for the multi-block decay time of price disparities and avoids the double-counting of the same opportunities.
6.1 Methodology
To prevent double-counting identical arbitrage opportunities that span several blocks, we adopt the methodology outlined in the empirical MAV assessments [13] and modify it for block-by-block analysis. It includes the following components:
  1. 1.
    Block Intervals: We analyze Binance ETH-USDC closing prices every second, and compare them with the spot price on the AMM-DEX at every block. The spot price on the DEX is determined based on the last swap within the block. The block production on Ethereum takes approximately 12 seconds, followed by around 2 seconds on Base, Optimism, and zkSync Era. On Arbitrum, blocks can be produced more quickly, approximately every 0.25 seconds. In such cases, we analyze the last block within the second. Figure 3(a) shows the ETH-USDC exchange rates on Binance and zkSync Era, highlighting the existence of price differences.
  2. 2.
    Price Re-Alignment Threshold: We only consider the price deviations that surpass a set threshold. This threshold is determined from the empirical distribution of price differences (see Fig. 3(b)), mainly focusing on outlier values (i.e., those beyond 1.5 times the interquartile range above the third quartile). This selected threshold ensures that prices are generally aligned but allows for adjustments to enhance flexibility.
  3. 3.
    Empirical MAV: Figure 3(c) illustrates the extent of price discrepancies. If a price difference continues for a prolonged duration of multiple blocks, we document only the highest MAV within each time segment of price discrepancy until re-alignment happens. This approach avoids counting the same ongoing price discrepancy multiple times.
  4. 4.
    Decay Time: Once all instances of empirical MAV (i.e., the peak MAV during periods of price misalignment) are identified, we determine the duration required for prices to re-align (i.e., when the magnitude of the price difference drops below the specified threshold).
((a)) Time series of prices for USDC-ETH.
((b)) Distribution of deltas.
((c)) Price differences at the block level, with the red dashed line indicating the threshold at which prices are assumed to have re-aligned. Missing bars represent blocks where no trading activity occurred on the AMM.
Figure 3: These plots illustrate the price differences for USDC-ETH between the AMM on zkSync and Binance. The AMM spot price fluctuates around the CEX price, showing how price misalignments first widen and subsequently narrow over a period of several minutes.
6.2 Empirical MAV from WETH-USDC pool
As stated previously, greater chances for arbitrage profit arise when there is a significant price disparity. Nevertheless, the possible profit is also dependent on the overall trade volume and the pool’s reserves.
Chain MAV % Volume Net LVR % Volume
Ethereum $52488479.88$ $0.15$ $126531170.12$ $0.36$
Arbitrum $2699032.60$ $0.04$ $20686037.49$ $0.28$
Arbitrum (USDC.e) $2707008.96$ $0.03$ $19332648.22$ $0.24$
Base $1293771.96$ $0.04$ $32663704.76$ $1.09$
Optimism (USDC.e) $690868.44$ $0.05$ $3893401.00$ $0.27$
zkSync Era (USDC.e) $20110.31$ $0.25$ $96016.54$ $1.20$
Table 4: Assessment of arbitrage opportunities within the WETH-USDC and WETH-USDC.e liquidity pools on Uniswap (v3) on Ethereum and its rollups from 31 December 2023 to 30 April 2024. The metrics encompass the total Maximal Arbitrage Value (MAV) and net Loss-Versus-Rebalancing (LVR), both in absolute terms and as percentages of trading volume.
Table 4 shows the total arbitrage opportunities (MAV) during the study periods, expressed as an absolute number and as a percentage of the trading volume. The greatest price differences occur on Ethereum, resulting in these arbitrage opportunities. However, these opportunities cannot be fully realized due to gas fees and MEV auction fees, which account for about 0.15% of the trading volume. The detected arbitrage opportunities on the Ethereum pool may remain untapped due to specific details in the Binance order book, which we cannot assess based on the 1-second data intervals.
The behavior seen in the rollups shows significant variations from those on Ethereum. In particular, MAV in the Arbitrum, Base, and Optimism pools ranges from 0.03% to 0.05% of trading volume, while in the zkSync pool, it is around 0.25%, suggesting that price discrepancies are not yet fully addressed in that rollup. Additionally, as expected, empirical MAV, which filters out identical arbitrage opportunities across multiple blocks, leads to lower rewards compared to cumulative net LVR. The LVR ranges from 0.24% to 1.2%, which is roughly five times higher than MAV.
Chain Points Avg MAV $V_{max}$ Avg Decay (s)
Ethereum $71135$ $737.87$ $499667.10$ 31
Arbitrum $137162$ $19.68$ $56476.35$ 6.9
Arbitrum (USDC.e) $164080$ $16.50$ $41370.11$ 8.8
Base $75729$ $17.08$ $22620.17$ 420
Optimism (USDC.e) $113428$ $6.09$ $8477.60$ 19
zkSync Era (USDC.e) $3970$ $5.07$ $1966.76$ 370
Table 5: Assessment of arbitrage opportunities within the WETH-USDC and WETH-USDC.e liquidity pools on Uniswap (v3) on Ethereum and its rollups from December 31st, 2023 to April 30th, 2024. The metrics encompass the quantity of identified arbitrage opportunities, mean Maximal Arbitrage Value (MAV), average transaction volume yielding maximal arbitrage, and mean decay time of the price disparities.
Table 5 provides further insight into the opportunities analyzed. The highest quantity of arbitrage opportunities is observed in the two Arbitrum pools, which can be attributed to higher block production rates. Base and Optimism also exhibit more arbitrage opportunities compared to the Ethereum pool. In particular, the average MAV is significantly lower on rollups, approximately 5–20 USD compared to over 700 USD on Ethereum. However, the transaction volume required to achieve this arbitrage is also lower (approximately 2.0002.0002.0002.0005.0005.0005.0005.000 USD on rollups and 0.5 million USD on Ethereum).
The average decay on rollups varies: for Arbitrum, it is approximately 7–9 seconds (with a 0.25-second block time), followed by 19 seconds on Optimism (2-second block time), 370 seconds on zkSync (1.05-second block time), and 420 seconds on Base (2-second block time). In each instance, the spans cover 10–20 blocks or more.
Figure 4: Empirical daily analysis of Maximal Arbitrage Value (MAV), price discrepancies, and liquidity in the current tick within the WETH-USDC liquidity pool on Uniswap (v3) deployed on the Arbitrum network.
Figure 4 illustrates the changes in price over time, the observed Maximal Arbitrage Value (MAV), and the pool’s virtual reserves within the current tick for the most active pool analyzed, WETH-USDC in Arbitrum. It specifically shows the maximum daily price difference, the daily total of observed MAV, and the end-of-day liquidity in the current pool. The data reveal that the daily MAV ranges from \$$2000$ to \$$140000$, with the lowest point occurring during low trading volumes between March and April 2024. The smaller price differences with Binance during this period result in a lower MAV. Importantly, the days with the highest daily MAV align with those with the largest price differences.
Figure 5 in the Appendix provides an analysis of liquidity pools on Ethereum, Base, Arbitrum, and Optimism. In particular, even with the introduction of native USDC on Arbitrum and its corresponding pool on Uniswap (v3), trading activity and associated arbitrage opportunities remain stable for both native USDC and bridged USDC tokens.
A significant correlation between arbitrage opportunities and the size of reserves and liquidity within the current tick is observed in the zkSync Era pool. The highest arbitrage opportunities occurred on days when liquidity within the current tick was at its maximum. In contrast, on Optimism, arbitrage opportunities are closely linked to price differences. This can be explained by the relatively small size of the pool on zkSync Era. Despite the rise in reserves on Base, there is no corresponding increase in arbitrage profits. This indicates that the estimation of arbitrage opportunities is complex and should be evaluated based on price discrepancies, fees, and pool liquidity.
7 Discussion
Arbitrage Presence.
Arbitrage reveals financial system inefficiencies. For AMMs on rollups, these may arise from security, infrastructure issues, or lack of understanding of rollup mechanics. Untapped arbitrage in rollups can be due to the absence of MEV auctions, low confidence in rollups’ upgradeability, and concerns about transaction finality. Misunderstandings about soft and hard finality at roll-up often lead to incorrect assumptions about bridging token durations. DeFi asset managers that focus on MEV auctions on Ethereum may avoid rollups due to limited trading volumes and liquidity, restricting arbitrage opportunities on roll-ups.
LVR as an Arbitrage Measure.
The main goal of the LVR metric is to compare the profitability between arbitrageurs and LPs rather than to measure arbitrage itself. In this research, we evaluate the absolute value of arbitrage opportunities, also known as net LVR (per block). As mentioned earlier, LVR assumes the immediate exploitation of arbitrage opportunities, and its direct application to rollups, where price discrepancies last across multiple blocks, would lead to double counting. There is an ongoing debate about whether arbitrageurs should be compared to LPs. LPs earn their rewards from fees on a pro-rata basis according to the liquidity they provide to the pool. In contrast, arbitrageurs are not assured of receiving arbitrage rewards, as they compete with each other to exploit the same opportunities. Therefore, liquidity provision and arbitrage are typically carried out by different asset management firms. Given the distinct infrastructure, software, and expertise needed, it is unlikely that a single asset management firm engages in both liquidity provision and arbitrage extraction.
MEV on Rollups.
As roll-up sequencers become decentralized, it can be anticipated that MEV auctions will be introduced to rollups. Currently, assuming trust in the centralized sequencer operators to prevent transaction reordering, MEV exploitation is limited to DEXs through back-running mechanisms within the smart contract. The back-running MEV can generate additional revenue for the DEX and, if redistributed, can also benefit the LPs contributing to the AMM liquidity pool. However, as previously mentioned, this arbitrage exploitation could be carried out by DEXs or specialized entities rather than LPs.
Positive and Negative MEV.
The ongoing discussion about the role of MEV, often referred to as MEV attacks due to their negative implications, continues. It is crucial to distinguish between front- and back-running MEVs, as well as their combined effects. Front-running MEV, especially when paired with back-running MEV, leads to various exploitative strategies, such as sandwich, imitation, and just-in-time liquidity attacks [32]. However, back-running MEV mainly supports arbitrage extraction and liquidation processes. Both arbitrage and liquidation play beneficial roles within DeFi ecosystems, fostering price balance and fair trading (arbitrage) and strengthening lending mechanisms (liquidations). As a result, there is a debate on whether rollup sequencers should allow only back-running MEV, as proposed by the time-boost mechanism [22].
Cross-Rollup Arbitrage.
Further research should explore cross-rollup arbitrage, especially considering the proposed use of shared sequencers across multiple roll-ups, which would enable atomic transactions between rollups [23]. Currently, arbitrage occurs mainly between AMMs and CEXs such as Binance, where trading volumes are highest, raising questions about the tokens to be exchanged across rollups. Arbitrage opportunities might emerge between Ethereum and rollups, particularly for tokens not listed on CEXs or with lower liquidity on CEXs. As trading shifts from Ethereum to roll-ups, with Ethereum increasingly acting as a settlement layer, cross-rollup arbitrage could become more common, especially when rollup liquidity surpasses that of CEXs and Ethereum’s AMMs.
8 Conclusions
This paper examines the DeFi dynamics in the Layer-2 ecosystem, focusing on AMMs and arbitrage on Ethereum rollups. We analyze swaps and arbitrage on the WETH-USDC pools on Uniswap (v3) on Ethereum, Arbitrum, Base, Optimism, and zkSync Era, noting significant user activity shifts from Ethereum to these rollups.
By analyzing the price discrepancies between Uniswap and Binance, we identified more than 0.5 million unexploited arbitrage opportunities on rollups. Significantly, these opportunities persisted on average for a period of 10 to 20 blocks. Thus, we revised the LVR metric to avoid double-counting arbitrage, recording only the highest arbitrage value (MAV) during sustained price differences over multiple blocks. MAV in Arbitrum, Base, and Optimism pools ranges from 0.03% to 0.05% of trading volume, while zkSync Era is around 0.25%, indicating unresolved price discrepancies. The empirical MAV approach, filtering repeated arbitrage, suggests that LVR metrics overestimated arbitrage by a factor of five.
We found that lower-volume swaps are more frequent on roll-ups compared to Ethereum. Swaps occur 2–3 times more often on rollups, but trade volumes are about five times lower. Block times impact rollup transactions; Ethereum averages one swap per block, while rollups have fewer swaps per block despite higher total swaps. Specifically, swaps occur every third block on Base, fifth on Optimism, and tenth on Arbitrum.
In conclusion, our research highlights unique behaviors and opportunities in the Layer-2 AMM-DEX ecosystem. Rollups offer not only lower gas fees, but also faster block production, leading to significant differences in the trading and arbitrage dynamics from that of Ethereum. These insights contribute to ongoing discussions on rollup design and optimization, particularly in the context of MEV auctions at sequencers.
Acknowledgements
This research article is a work of scholarship and reflects the authors’ own views and opinions. It does not necessarily reflect the views or opinions of any other person or organization, including the authors’ employer. Readers should not rely on this article for making strategic or commercial decisions, and the authors are not responsible for any losses that may result from such use. One of the authors’ research was supported by the EPSRC CDT in Mathematics of Random Systems (EPSRC Grant EP/S023925/1).
Appendix 0.A Appendix
((a)) Ethereum
((b)) Base
((c)) Arbitrum
((d)) Arbitrum (USDC.e)
((e)) Optimism (USDC.e)
((f)) zkSync Era (USDC.e)
Figure 5: Empirical Maximal Extractable Value (MAV), price misalignment with Binance, and liquidity within the spot price tick of the WETH-USDC pool on Uniswap (v3) across Ethereum and its rollups from December 31st, 2023, to April 30th, 2024.

### v2

@@
-Layer-2 Arbitrage: An Empirical Analysis of Swap Dynamics and Price Disparities on Rollups
+Cross-Rollup MEV:
 Abstract
-This paper explores the dynamics of Decentralized Finance (DeFi) within the Layer-2 ecosystem, focusing on Automated Market Makers (AMM) and arbitrage on Ethereum rollups.
-We observe significant shifts in trading activity from Ethereum to rollups, with swaps on rollups happening 2–3 times more often, though, with lower trade volume.
-By examining the price differences between AMMs and centralized exchanges, we discover over 0.5 million unexploited arbitrage opportunities on rollups.
-Remarkably, we observe that these opportunities last, on average, 10 to 20 blocks, requiring adjustments to the LVR metrics to avoid double-counting arbitrage.
-Our results show that arbitrage in Arbitrum, Base, and Optimism pools ranges from 0.03% to 0.05% of trading volume, while in zkSync Era it oscillates around 0.25%, with the LVR metric overestimating arbitrage by a factor of five.
-Rollups offer not only lower gas fees, but also provide faster block production, leading to significant differences compared to the trading and arbitrage dynamics of Ethereum.
-Keywords: Arbitrage, Automated Market Maker, Rollup, MEV
+This study quantifies the potential non-atomic MEV on Layer-2 (L2) blockchains by measuring the arbitrage opportunities between cross-rollup and DEX-CEX.
+Over recent years, we observe a shift in trading activities from Ethereum to rollups, with swaps on rollups occurring 2–3 times more frequently, albeit with lower trade volumes.
+By analyzing the costs of swap on L2s and price discrepancies cross-rollup and DEX-CEX, we identify more than $500000$ unexplored arbitrage opportunities.
+In particular, we find that these opportunities persist, on average, for 10 to 20 blocks, necessitating the modification of the Loss Versus Rebalancing (LVR) metric to prevent double-counting.
+Our findings indicate that the arbitrage opportunities in Arbitrum, Base, and Optimism range between 0.03% and 0.05% of the trading volume, while in the ZKsync Era it fluctuates around 0.25%.
 1 Introduction
-Ethereum has undergone a series of substantial enhancements aimed at increasing its functionality, addressing scalability challenges, and improving user interactions [2].
-A pivotal upgrade was the transition from Proof of Work (PoW) to Proof of Stake (PoS), referred to as The Merge, which occurred throughout 2022 and 2023.
-After that, the most significant milestone was the Dencun upgrade, implemented on March 13th, 2024.
-It introduced Proto-Danksharding, specifically designed for rollups: Layer-2 (L2) scaling solutions.
-This upgrade marked the initial phase of Ethereum’s rollup-centric roadmap and introduced blobs, temporary storage for rollup transaction data, streamlining verification, and reducing data storage costs [2].
-Consequently, the gas fees on rollups have been significantly reduced, as shown in Figure 1.
-Figure 1: Median transaction fees, measured in Gwei, for selected Ethereum rollups, pre- and post-Dencun Upgrade on March 13th, 2024.
-The reduction in gas fees has influenced the use of Decentralized Finance (DeFi) on rollups.
-In recent years, the Total Value Locked (TVL) and the trading volume on rollups have exhibited a notable increase, with certain rollups now surpassing Ethereum in daily transaction numbers [1, 3].
-Furthermore, the dynamics of L2 DeFi is not solely altered by reduced gas fees.
-The second important factor is the faster block production, typically less than 2 seconds in contrast to Ethereum’s 12 seconds [28].
-The combination of lower gas fees and faster block production enables arbitrageurs to capitalize on smaller price discrepancies between trading venues more quickly.
-Although arbitrage transactions are performed predominantly on Ethereum through Maximal Extractable Value (MEV) Boost Auctions [16], an equivalent MEV boost mechanism is not present in rollups.
-The absence of MEV boost auctions on rollups can be explained by the fact that most rollups rely on centralized sequencers to aggregate transactions into L1 batches.
-The centralized sequencers and the associated censorship risk remain a concern within the L2 DeFi ecosystem.
-Additional concerns encompass transaction finality, the time period until an L2 transaction becomes irrevocable from the rollup (soft finality) or the underlying layer-1 blockchain (hard finality) [31].
+Decentralized Finance (DeFi) is transitioning towards rollups, the Layer-2 (L2) scaling solutions for Ethereum.
+The primary reason is that rollups provide equivalent security guarantees as Ethereum while significantly reducing gas fees [41].
+Rollups offload complex computations outside the Ethereum network, storing only compressed batches of transactions on-chain.
+In March 2023, Ethereum underwent the Dencun upgrade [13], the most substantial change since the Merge, which further lowered gas fees on rollups, as shown in Figure 1.
+The Dencun upgrade introduced Proto-Danksharding and blobs — temporary data storage specifically designed to optimize rollups performance [14].
+[…]
+The trading volume on Decentralized Exchanges (DEXs) within L2s is increasing, with certain rollups surpassing Ethereum’s daily swap counts [11, 30].
+Then, DEXs within L2s present new opportunities for arbitrageurs aiming to capitalize on price disparities across rollups or between rollups and Centralized Exchange (CEX).
+However, while arbitrage on Ethereum is currently facilitated through Maximal Extractable Value (MEV) boost auctions [24], such mechanisms are yet unavailable in rollups.
+Most L2s rely on centralized sequencers to produce blocks and aggregate them into Layer-1 batches, employing the first-come-first-served rule.
+In addition, block production on L2s usually takes less than 2 seconds, as opposed to Ethereum’s 12-second block time [41].
+The combination of reduced gas costs and faster block production allows arbitrageurs to exploit minor price differences at a faster pace.
+Arbitrage can be classified into two types: i) atomic arbitrage within DEXs on the same blockchain (cycling arbitrage [43]) or ii) non-atomic arbitrage involving CEX or DEX on a different blockchain [24].
+Currently, 71% of the MEV transaction volume on Ethereum is related to arbitrage, and in excess of 60% of that arbitrage volume comes from nonatomic transactions [12, 16, 24].
+Cycling arbitrage on L2 can be executed as atomic transactions [43], but non-atomic arbitrage faces a risk of state (price) drift.
+With the significantly reduced gas fees following the Dencun upgrade (see Figure 1), non-atomic arbitrage can be exploited by posting a large amount of prospective arbitrage transactions that are reverted if unsuccessful.
+The number of revered transactions on L2s increased from 2% to 6% after the Dencun upgrade [15].
+Another strategy used by L2 arbitrageurs is the physical location of centralized sequencers, a service sometimes provided by DeFi solvers [9].
+Transactions originating from the same data center as the centralized sequencer are likely to be executed first due to shorter network delays [10, 32, 44].
+However, risks associated with centralized sequencers related to censorship remain a concern within L2 DeFi.
+Additional risks include the transaction finality, the time period until an L2 transaction becomes irrevocable from the rollup (soft finality) or the underlying L1 blockchain (hard finality) [47].
 Rollups can be categorized into Optimistic and Zero-Knowledge (ZK) rollups based on transaction validity and proof mechanisms [19].
@@
 However, ZK-rollups offer higher transaction compression, throughput, and faster hard finality, ranging from minutes to hours.
-The aforementioned factors — reduced gas fees, absence of MEV auctions, varying finality speeds, and accelerated block production — contribute to distinct DeFi dynamics on rollups, which warrant further investigation.
-With the emergence of decentralized rollup sequencers [26], there arises a potential risk that adverse practices associated with MEV on Ethereum [27] transfer to the L2 DeFi ecosystem.
+The aforementioned factors — reduced gas fees, absence of MEV auctions, varying finality times, and faster block production — contribute to distinct arbitrage dynamics on rollups, which requires further investigation.
+With the emergence of decentralized sequencers [38], the adverse practices associated with front-running MEV on Ethereum [10, 40, 42] transfer to L2 DeFi.
 To address this concern, a time-boost mechanism within sequencers has been proposed [34].
-This mechanism is specifically designed to allow back-running MEVs while preventing front-running activities.
-Front-running MEV can facilitate various forms of adversarial actions, such as sandwich attacks on traders [19] and Just-In-Time (JIT) liquidity attacks on Liquidity Providers (LPs) within AMM-DEXs [32].
-In contrast, back-running MEV enables immediate arbitrage at DEXs or liquidations at lending protocols.
+This mechanism is specifically designed to allow only back-running MEVs while preventing front-running.
+Front-running MEV can enable various forms of adversarial actions, such as sandwich attacks on traders [28, 36] and Just-In-Time (JIT) liquidity attacks on Liquidity Providers (LPs) [48].
+In contrast, back-running MEV enables arbitrage in DEXs and liquidations at lending protocols.
 Faster arbitrage execution benefits traders, as it eliminates price discrepancies between trading venues.
 Additionally, MEV time-boost auctions can generate new revenue for sequencer operators.
-This paper contributes to the research on L2 DeFi by empirically examining the differential swap dynamics on rollups and Ethereum, considering the factors mentioned above, such as lower gas fees, faster block production, and the absence of MEV auctions.
-Furthermore, we quantify the potential arbitrage opportunities across rollups and between rollups and CEXs that can be the basis for the revenue estimation from the time-boost MEV auctions in the rollup sequencers.
-Related Work.
+1.1 Contributions
+This work contributes to the research on DeFi and MEV on L2s by quantifying the potential of non-atomic arbitrage on rollups.
+We conduct an empirical investigation into price disparities cross-rollups and measure the decay time.
+Our findings advance ongoing research on the applications of shared sequencer and time-boost (MEV) auctions.
+Our key contributions are as follows.
+L2 Non-atomic Arbitrage and Its Decay Over Time.
+By examining the price differences between Binance and Uniswap (v3) WETH-USDC pools on L2s, we discover over 0.5 million unexploited arbitrage opportunities on rollups.
+The Maximal Arbitrage Value (MAV) [21] on Arbitrum, Base, and Optimism pools varies from 0.03% to 0.05% of trading volume, whereas in ZKsync Era, it is approximately 0.25%.
+We show that episodes of price misalignment often last for 10–20 blocks in rollups; we further present a framework to capture only one MAV during each misalignment period, avoiding double-counting.
+The empirical MAV method that filters out repeated arbitrage results in 5×5$\times$5 × lower arbitrage profits compared to the Loss-Versus-Rebalancing (LVR) metric.
+Cross-Rollup MEV.
+We estimate the potential for cross-rollup MEV by analyzing the price disparities among WETH-USDC Uniswap (v3) pools on major L2s.
+The arbitrageurs profits are influenced by the price impact (of AMM) and, thus, are the highest between the AMM pools with the highest token reserves.
+2 Related Work
 There exists a substantial body of research on DeFi, particularly focused on AMM-DEXs.
@@
 [46] and empirical investigations of Uniswap (v3) on the Ethereum blockchain [17, 20, 26, 27, 33] are prominent.
-Most of these studies evaluate the profitability of LPs, specifically addressing impermanent loss.
+Most of these studies evaluate the profitability of Liquidity Providers (LPs), specifically addressing impermanent loss.
 The analysis by Adams et al.
-[4] delves into the costs and dynamics of swaps on Uniswap (v3) on Ethereum.
+[1] analyzes the costs and dynamics of swaps on Uniswap (v3) on Ethereum.
 Our research extends this by breaking down swap fees on rollups and assessing the influence of blobs and block time on swap dynamics.
@@
 The Loss Versus Rebalancing metric (LVR), which compares arbitrageur profits with those of LPs, was introduced by Milionis et al.
-[24, 25] and empirically measured by Fritsch et al.
+[37] and empirically measured by Fritsch et al.
 [18].
@@
 First, LVR assumes that the price discrepancies between AMMs and target exchanges, such as CEXs, disappear after each block.
-Our findings confirm this assumption for Uniswap (v3) on Ethereum; nonetheless, the decay period of price discrepancies on rollups extends over multiple blocks.
+Our findings confirm this assumption for Uniswap (v3) on Ethereum; however, the decay period of price discrepancies on rollups extends over multiple blocks.
 Our methodology adjusts the LVR to this scenario and prevents double counting of the same arbitrage opportunities.
@@
 [21].
-Our paper builds on that research by providing: i) a comprehensive analysis of rollups (Arbitrum, Base, Optimism, in addition to zkSync Era) alongside Ethereum, ii) enhanced precision in market price accuracy and per-block analysis, (improving from a 1-minute to a 1-second resolution), and iii) an in-depth examination of Uniswap (v3), focusing on the most actively traded pools.
+Our paper builds on that research by providing: i) a comprehensive analysis of rollups (Arbitrum, Base, Optimism, in addition to ZKsync Era) alongside Ethereum, ii) enhanced precision in market price accuracy and per-block analysis, improving from a 1-minute to a 1-second resolution, and iii) an in-depth examination of Uniswap (v3), focusing on the most actively traded pools.
 Heimbach et al.
@@
 Similarly, we establish the minimum threshold for arbitrage opportunities and derive the arbitrage value using AMM formulas.
-However, this work seeks to quantify the unexploited arbitrage, which occasionally persisted for several minutes, in contrast to the arbitrage realized via MEV on Ethereum.
+However, our work seeks to quantify the unexploited arbitrage, which occasionally persisted for several minutes, in contrast to the arbitrage realized via MEV on Ethereum.
 Other studies on arbitrage [4, 5, 31] focus primarily on AMM-DEX on Ethereum and/or cyclic arbitrage opportunities.
 Torres et al.
-[29] is the first effort to quantify MEV in roll-ups.
+[43] is the first effort to quantify MEV in rollups.
 Despite the absence of MEV auctions, this research examines the heuristic for back-running arbitrage MEV, specifically focusing on swap events occurring within the same transaction in the block, within the context of cylindrical arbitrage.
 Our objective is to assess the potential for non-atomic arbitrage, involving CEXs, or cross-rollup price discrepancies.
-Mamageishvili et al.
+Finally, Mamageishvili et al.
 [34] introduce a time boost mechanism within rollup sequencers.
@@
 However, these studies lack empirical quantification of arbitrage opportunities within rollups, a gap that this work seeks to address.
-Contributions.
-This paper advances DeFi research within rollups.
-We conducted an empirical investigation into the distinct swap dynamics in optimistic and ZK rollups, as compared to Ethereum, taking into account specific rollup characteristics, such as lower gas fees, faster block production, and the absence of MEV auctions.
-Subsequently, we analyze the arbitrage opportunities on rollups that can be exploited by arbitrageurs, thereby providing new revenue streams for sequencers offering back-running MEV auctions.
-Our principal contributions are as follows.
-  • Examination of Arbitrage and Swapping Costs on Rollups: We conduct a comparative study of the costs of swapping and capturing arbitrage on rollups compared to Ethereum.
-Our findings indicate that, unlike Ethereum, transactions on rollups occur 2–3 times more frequently but are smaller in size.
-Furthermore, swaps on rollups are generally not impacted by front- or back-running block slippage typically caused by MEV.
-Block times impact rollup transactions; Ethereum averages one swap per block, while rollups have fewer swaps per block despite higher total swaps.
-Specifically, swaps occur every third block on Base, fifth on Optimism, and tenth on Arbitrum.
-  • Decay of Price Disparities: We show that price differences last for 10–20 blocks in rollups, necessitating modifications in the empirical assessment of LVR metrics.
-We present a framework for empirically estimating arbitrage in cases where the decay period of price differences spans multiple blocks, capturing only the maximum arbitrage value (MAV) during prolonged price disparities.
-The empirical MAV method, which filters out repeated arbitrage, indicates that LVR metrics have overestimated arbitrage by a factor of five.
-  • Estimation of Arbitrage on Rollups: By examining the price differences between Uniswap (v3) and Binance, we discovered over 0.5 million unexploited arbitrage opportunities on rollups.
-The MAV metric on Arbitrum, Base, and Optimism pools varies from 0.03% to 0.05% of trading volume, whereas in zkSync Era, it is approximately 0.25%, highlighting significant unresolved price differences.
-Paper Organization.
-Section 2 offers foundational knowledge on rollups and AMMs.
-Section 3 details the dataset utilized in our analysis, while Section 4 breaks down swap costs on Ethereum rollups, both pre- and post-Decun (blobs).
-Section 5 sets the theoretical foundation for arbitrage in AMMs and its estimation framework in rollups.
-Section 6 presents the results of our empirical investigations.
-Section 7 examines the significance of arbitrage in rollups and proposes future research directions.
-Lastly, Section 8 summarizes our findings as well as some final remarks.
 3 Background
-This section introduces rollups, a new mechanism of layer-2 blockchain scaling, and formalizes the blockchain terminology used throughout the manuscript.
-Furthermore, it presents the basics of Decentralized Exchanges (DEXs) and Automated Market Makers (AMMs).
-2.1 Layer-2 Blockchains
 There are two main approaches — Layer-1 (L1) and Layer-2 (L2) scaling — to tackle blockchain scalability challenges and the associated increases in gas fees.
 L1 scaling involves creating a new blockchain with its own consensus mechanisms and a unique physical infrastructure to support the network.
-In contrast, L2 scaling uses a different method: it carries out computations off the main blockchain while recording their results on the underlying chain.
+In contrast, L2 scaling uses a different method: it carries out computations off the main blockchain while recording their results on the underlying L1 chain.
 The main L2 scaling solutions include state (payment) channels, plasma, validium, and rollups.
@@
 Rollup.
-A rollup functions as a virtual blockchain, producing transaction blocks and recording them on the main chain as compressed batches.
-However, it operates with the risk of adversarial actions such as stopping block production, generating invalid blocks, withholding data, or engaging in other malicious activities.
+A rollup functions as a virtual blockchain, producing transaction and blocks and recording them on the main chain as compressed batches.
+However, it operates with the risk of adversarial actions such as stopping block production, generating invalid blocks, withholding data, or engaging in other malicious activities [7].
 To ensure the correctness of the transaction state in these untrusted environments, rollups use one of two approaches: the optimistic approach and zero-knowledge (ZK) proofs.
-Figure 2 illustrates the rollup architecture, highlighting its principal components: sequencers and verifiers.
+Figure 5, in §0.A, illustrates the rollup architecture, highlighting its principal components: sequencers and verifiers.
 Sequencers consolidate transactions to be rolled up onto the L1 chain, thereby achieving gas fee reductions.
 Verifiers, which are smart contracts on L1, validate the transaction block created by the sequencer, thus ensuring its correctness.
-Furthermore, ZK-rollups utilize provers responsible for generating ZK validity proofs.
-Figure 2: Architecture of a rollup
+Furthermore, ZK-rollups utilize provers responsible for generating ZK validity proofs [8].
 Soft and Hard Finality.
@@
 Dencun Upgrade and Blobs.
-On 13 March 2024, Ethereum underwent the Dencun Upgrade, which introduced temporary data storage known as “blobs” [2].
-Blobs are designed to enhance the efficiency of Ethereum rollups.
-Ethereum nodes store blob data for 18 days, during which rollups validate their transactions.
-After this 18-day period, the blobs are removed from the Ethereum blockchain.
-Rollups are not required to keep transaction details from blobs; however, they typically store off-chain all transaction data for a better user experience.
-2.2 Automated Market Makers
+On March 13, 2024, Ethereum implemented the Dencun upgrade, introducing temporary data storage called “blobs” [13].
+Blobs improve the efficiency of Ethereum rollups by storing data for 18 days, during which rollups validate transactions.
+After this period, the blobs are deleted from the blockchain.
+While rollups aren’t required to retain blob transaction data, they usually store it off-chain to provide users with a complete transaction history.
+4 Model for Arbitrage
+Price disparities between different trading platforms (both DEXs and CEXs) present certain arbitrage opportunities.
+In this section, we derive the formula for the trading volume that equals prices and maximizes the arbitrageur’s profit.
+This value is referred to as the net LVR [37], or the Maximal Arbitrage Value (MAV) [21].
+For more details on AMMs, we refer the reader to Appendix §0.A.
+Maximal Arbitrage Value (MAV).
+Consider a scenario of two tokens - X and Y - in which there is a price discrepancy between a CEX and an AMM, the CEX having a substantially higher market depth, resulting in zero market impact.
+Let $P_{CEX}$ and $P_{AMM}$ be the exchange prices of the token X for Y at CEX and AMM, respectively, with $g$ being the CEX fee and $f$ LP fee at AMM.
+When $P_{AMM}$ < $P_{CEX}$, the arbitrageur buys $\Delta x(1-f)$ tokens X at AMM and sells them at CEX. His profit, denominated in Y, is
+$V(\Delta x)=\Delta x(1-f)P_{CEX}(1-g)-\Delta xP_{AMM}\rho(\Delta x)$ (1)
+where $\rho(\Delta x)$ is the percentage price impact of his transaction on the AMM price.
+Assuming that $P_{CEX}$ < $P_{AMM}$, the arbitrageur buys $\Delta x(1-g)$ tokens X at CEX for $\Delta xP_{CEX}$ and sells them at AMM. His profit, denominated in Y, is
+$V(\Delta x)=\Delta x(1-g)P_{AMM}(1-\rho(\Delta x(1-g)))(1-f)-\Delta xP_{CEX}$ (2)
+Arbitrage between a CPMM and CEX.
+Let us assume that the respective reserves of tokens X and Y in the AMM liquidity pool are $x$ and $y$, and $P_{AMM}$ < $P_{CEX}$.
+Assuming that AMM follows CPMM, for a swap of $\Delta x$ for $\Delta y$, the percentage price impact is given by $\rho(\Delta x)=\frac{\Delta y}{y}$ [21, 46].
+As LP fee $f\Delta x$ is not part of token reservers, then $xy=(x-\Delta x)(y+\Delta y)$ and $1+\frac{\Delta y}{y}=\frac{1}{1-\frac{\Delta x}{x}}$.
+Thus, the MAV can be expressed as:
+$V(\Delta x)=\Delta x\left[(1-f)(1-g)P_{CEX}-\frac{P_{AMM}}{1-\frac{\Delta x}{x}}\right]$ (3)
+We can compute the first derivative of Eq. (3) with respect to $\Delta x$ to find the max points, receiving:
+$\Delta x_{max}=x\cdot(1-\frac{\sqrt{P_{AMM}}}{\sqrt{(1-f)(1-g)P_{CEX}}})$ (4)
+And, by substituting (6) into (3) the MAV is written as:
+$V(\Delta x_{max})=x\cdot(1-f)\left[\sqrt{(1-g)P_{CEX}}-\sqrt{\frac{P_{AMM}}{1-f}}\right]^{2}$ (5)
+Analogously, we can find MAV when $P_{AMM}>P_{CEX}$:
+$\Delta x_{max}=\frac{x}{1-g}\cdot(\frac{\sqrt{(1-f)(1-g)P_{AMM}}}{\sqrt{P_{CEX}}}-1)$ (6)
+$V(\Delta x_{max})=x\cdot\left[\sqrt{(1-f)P_{AMM}}-\sqrt{\frac{P_{CEX}}{1-g}}\right]^{2}$ (7)
+Probability of Successful Transaction.
+Let us assume that $P_{CEX}=P_{AMM}(1+\epsilon)$, where $\epsilon>0$ and the probability of successful execution of the arbitrage transaction of volume $\Delta x$ is $\mathbb{P}(\Delta x,\epsilon)$.
+Then, the payoff of the arbitrageour is
+$\hat{V}(\Delta x,\epsilon)=V(\Delta x,\epsilon)\cdot\mathbb{P}(\Delta x,\epsilon)-gasfee$ (8)
+$\hat{V}(\Delta x,\epsilon)=P_{AMM}\Delta x(1+\epsilon-\frac{x}{x-\Delta x})\cdot\mathbb{P}(\Delta x,\epsilon)-gasfee$
+Arbitrage between two CPMMs.
+Following analogous calculation, we can derive the formula for $x_{max}$ and MAV between two CPMMs.
+Assuming reserves of token X in their liquidity pools given by $x_{1}$ and $x_{2}$, and spot prices $P_{1}$ and $P_{2}$:
+$\Delta x_{max}=\frac{\sqrt{P_{1}}-\sqrt{P_{2}}}{\frac{\sqrt{P_{1}}}{x_{1}}-\frac{\sqrt{P_{2}}}{x_{2}}}$ (9)
+5 Data Collection
+We analyze swap data from Ethereum and its rollups with the highest trading volumes.
+This includes Arbitrum, Base, and Optimism (optimistic rollups), and ZKsync Era (ZK-rollup).
+For each blockchain, we focus on the most actively traded pool, WETH-USDC on Uniswap (v3), during the period from December 31st, 2023 and April 30th, 2024.
+The data set is sourced from the blockchain archive nodes provided by Nansen [39].
+The number of analyzed swaps, transactions, and block ranges is detailed in Table 1.
+Using the event logs of the Swap method, we recalculate the historical spot prices and liquidity in ticks within the USDC-WETH pools.
+For Uniswap (v3), the spot price after the swap is obtained from sqrtPriceX96 in the event logs.
+Market data for the ETH-USDC exchange rate on CEX are sourced from Binance APIs [6].
+We examine 1-second Binance closing prices, noting that block production on Ethereum occurs every 12 seconds, while it is around 1–2 seconds on Base, Optimism, and ZKsync Era.
+On Arbitrum, block production can be as fast as 0.25 seconds.
+Thus, for each blockchain, except for Arbitrum, computations are performed at the block level.
+For Arbitrum, we analyze the last block in a second.
+The DEX spot price is determined by the last swap within the block.
+Chain Swaps Transactions Blocks Block Range
+Ethereum $761005$ $749818$ $475409$ 1890889618908896189088961890889619771559197715591977155919771559
+Arbitrum $2400000$ $2367361$ $1709619$ 187373628187373628187373628187373628206540031206540031206540031206540031
+Arbitrum (USDC.e) $2258469$ $2232543$ $1648839$ 165788868165788868165788868165788868206540037206540037206540037206540037
+Base $1687530$ $1636196$ $1145789$ 863892986389298638929863892913866123138661231386612313866123
+Optimism (USDC.e) $1186780$ $1136839$ $902789$ 114234215114234215114234215114234215119461410119461410119461410119461410
+ZKsync Era (USDC.e) $46417$ $46379$ $45364$ 2290992322909923229099232290992332843035328430353284303532843035
+Table 1: List of analyzed WETH-USDC and WETH-USDC.e liquidity pools on Uniswap (v3) on Ethereum and its rollups, spanning the period from 31st of December 2023 to 30th of April 2024.
+Native USDC and Bridged USDC.e.
+USDC is a fiat-backed stablecoin issued by Circle.
+It takes one of two forms: native or bridged token.
+The native USDC is directly backed by Circle’s off-chain reserves, while USDC.e is bridged, typically from Ethereum.
+The bridged stablecoin, USDC.e, is faster to launch on new blockchains, while the issuance of the native USDC stablecoin often follows.
+As a result, in most rollups, USDC and USDC.e tokens coexist for legacy purposes.
+Dynamics and Cost of Swapping on L2s.
+We break down these costs and present the characteristics of WETH-USDC swaps on rollups on Tables 4 and 5 in §0.B, while Table 6 (also in §0.B) breaks down the costs of swapping.
+The corresponding values of Ethereum are provided as a benchmark.
+As shown in Table 4, swap transactions on rollups have a lower volume compared to those on Ethereum, although they occur more frequently.
+On average, there are 2–3 times more swaps on rollups, but the trading volume is about five times lower.
+Consequently, the median swap volume on Ethereum is $4239.45$ USD, while it is $2201.23$ USD on Arbitrum, $173.46$ USD on Base, $331.28$ USD on Optimism, and only $20$ USD on ZKsync Era; see Appendix §0.B for details.
+6 Empirical Results
+((a)) Time series of prices for USDC-ETH.
+((b)) Distribution of deltas.
+((c)) Price differences at the block level, with the red dashed line indicating the threshold at which prices are assumed to have re-aligned.
+Missing bars represent blocks where no trading activity occurred on the AMM.
+Figure 2: These plots illustrate the price differences for USDC-ETH between the AMM on ZKsync and Binance.
+The AMM spot price fluctuates around the CEX price, showing how price misalignments first widen and subsequently narrow over a period of several minutes.
+When price discrepancies arise between trading platforms, arbitrageurs can take advantage of these opportunities to profit from price corrections.
+This activity benefits both the arbitrageurs, who gain from the price differences, and traders, who enjoy promptly adjusted and accurate exchange rates.
+This section examines the WETH-USDC price misalignment.
+We measure arbitrage opportunities between rollup AMMs and CEX, and across rollups.
+We also analyze the decay time of these discrepancies and quantify the arbitrage opportunities using the Maximal Arbitrage Value (MAV) approach.
+6.1 Methodology
+To avoid double-counting the same arbitrage opportunities that span multiple blocks, we use the MAV methodology from Gogol et al[21] and adapt it for block-by-block analysis. It includes:
+Block intervals.
+We analyze the Binance ETH-USDC closing prices at every second and compare them with the spot price on the AMM-DEX at every block.
+The spot price on the DEX is determined based on the last swap within the block.
+The block production on Ethereum takes approximately 12 seconds, followed by around 2 seconds on Base, Optimism, and ZKsync Era.
+On Arbitrum, blocks can be produced more quickly, approximately every 0.25 seconds.
+In such cases, we analyze the last block within the second.
+Figure 2(a) shows the ETH-USDC exchange rates on Binance and ZKsync Era, highlighting the existence of price differences.
+Price re-alignment threshold.
+We only consider the price deviations that surpass a set threshold.
+This threshold is determined from the empirical distribution of price differences (see Figure 2(b)), mainly focusing on outlier values (i.e., those beyond 1.5 times the interquartile range above the third quartile).
+This selected threshold ensures that prices are generally aligned but allows for adjustments to enhance flexibility.
+Empirical MAV.
+Figure 2(c) illustrates the extent of price discrepancies.
+If a price difference continues for a prolonged duration of multiple blocks, we document only the highest MAV within each time segment of price discrepancy until re-alignment happens.
+This approach avoids counting the same ongoing price discrepancy multiple times.
+Decay time.
+Once all instances of empirical MAV (i.e., the peak MAV during periods of price misalignment) are identified, we determine the duration required for prices to re-align (i.e., when the magnitude of the price difference drops below the specified threshold).
+6.2 Empirical MAV from WETH-USDC Pool
+Chain MAV % Volume Net LVR % Volume
+Ethereum $52488479.88$ $0.15$ $126531170.12$ $0.36$
+Arbitrum $2699032.60$ $0.04$ $20686037.49$ $0.28$
+Arbitrum (USDC.e) $2707008.96$ $0.03$ $19332648.22$ $0.24$
+Base $1293771.96$ $0.04$ $32663704.76$ $1.09$
+Optimism (USDC.e) $690868.44$ $0.05$ $3893401.00$ $0.27$
+ZKsync Era (USDC.e) $20110.31$ $0.25$ $96016.54$ $1.20$
+Table 2: Assessment of arbitrage opportunities within the WETH-USDC and WETH-USDC.e liquidity pools on Uniswap (v3) on Ethereum and its rollups from 31st of December 2023 to the 30th of April 2024.
+The metrics encompass the total Maximal Arbitrage Value (MAV) and net Loss-Versus-Rebalancing (LVR), both in absolute terms and as percentages of trading volume.
+Arbitrage profit opportunities increase with price disparity, and potential profit depends on trade volume and pool reserves.
+Table 2 shows the total arbitrage opportunities (MAV) during the study periods, as an absolute number and a percentage of trading volume.
+Although the highest price differences are on Ethereum, offering about 0.15% of trading volume in arbitrage opportunities, these may not be fully realized due to gas fees and MEV auction costs.
+Additionally, specific details in the Binance order book, unavailable from 1-second data interval research, can contribute to untapped opportunities.
+Non-atomic arbitrage on rollups presents significant differences from Ethereum.
+MAV in Arbitrum, Base, and Optimism pools ranges from 0.03% to 0.05% of trading volume, whereas in ZKsync, it is around 0.25%, resulting from price misaligment occurring in multiple blocks.
+Empirical MAV, filtering identical arbitrage opportunities across blocks, results in lower rewards than cumulative net LVR.
+LVR ranges from 0.24% to 1.2%, about five times higher than MAV.
+Chain Points Avg MAV $V_{max}$ Avg Decay (s)
+Ethereum $71135$ $737.87$ $499667.10$ 31
+Arbitrum $137162$ $19.68$ $56476.35$ 6.9
+Arbitrum (USDC.e) $164080$ $16.50$ $41370.11$ 8.8
+Base $75729$ $17.08$ $22620.17$ 420
+Optimism (USDC.e) $113428$ $6.09$ $8477.60$ 19
+ZKsync Era (USDC.e) $3970$ $5.07$ $1966.76$ 370
+Table 3: Arbitrage opportunities within the WETH-USDC and WETH-USDC.e liquidity pools on Uniswap (v3) on Ethereum and its rollups from Dec.
+31st, 2023 to Apr.
+30th, 2024.
+The metrics include the quantity of identified arbitrage opportunities, average Maximal Arbitrage Value (MAV), average transaction volume yielding maximal arbitrage, and mean decay time of the price disparities.
+Table 3 provides further insight into the opportunities analyzed.
+The highest quantity of arbitrage opportunities is observed in the two Arbitrum pools, which can be attributed to higher block production rates.
+Base and Optimism also exhibit more arbitrage opportunities compared to the Ethereum pool.
+In particular, the average MAV is significantly lower on rollups, approximately 5–20 USD compared to over 700 USD on Ethereum.
+However, the transaction volume required to achieve this arbitrage is also lower (approximately 2.0002.0002.0002.0005.0005.0005.0005.000 USD on rollups and 0.5 million USD on Ethereum).
+Additionally, the average decay on rollups varies: for Arbitrum, it is approximately 7–9 seconds (with a 0.25-second block time), followed by 19 seconds on Optimism (2-second block time), 370 seconds on ZKsync (1.05-second block time), and 420 seconds on Base (2-second block time).
+In each case, the spans cover 10–20 blocks or more.
+Figure 3: Empirical daily analysis of Maximal Arbitrage Value (MAV), price discrepancies, and liquidity in the current tick within the WETH-USDC liquidity pool on Uniswap (v3) deployed on the Arbitrum network.
+Figure 3 illustrates the changes in price over time, the MAV, and the pool’s virtual reserves within the current tick for the most active pool analyzed, WETH-USDC in Arbitrum.
+It specifically shows the maximum daily price difference, the daily total of observed MAV, and the end-of-day liquidity in the current pool.
+The data reveal that the daily MAV ranges from \$$2000$ to \$$140000$, with the lowest point occurring during low trading volumes between March and April 2024.
+The smaller price differences with Binance during this period result in a lower MAV.
+Importantly, the days with the highest daily MAV align with those with the largest price differences.
+Figure 6 in §0.C provides an analysis of MAV on Ethereum, Base, Arbitrum, and Optimism.
+In particular, even with the introduction of native USDC on Arbitrum and its corresponding pool on Uniswap (v3), trading activity and associated arbitrage opportunities remain stable for both native USDC and bridged USDC tokens.
+A significant link between arbitrage opportunities and the size of reserves and liquidity within the current tick is observed in the ZKsync Era pool.
+The highest arbitrage opportunities occurred on days when liquidity within the current tick was at its maximum.
+In contrast, on Optimism, arbitrage opportunities are closely linked to price differences.
+This can be explained by the relatively small size of the pool on ZKsync Era.
+Despite the rise in reserves on Base, there is no corresponding increase in arbitrage profits.
+This indicates that the estimation of arbitrage opportunities is complex and should be evaluated based on price discrepancies, fees, and pool liquidity.
+6.3 Cross-Rollup Arbitrage
+Cross-rollup arbitrage involves carefully planned steps on at least two rollups and revenue from arbitrage must cover all fees.
+If an token’s price is lower on one rollup than another, arbitrageurs buy low on the first rollup and sell high on the second to profit from the price difference.
+We show the WETH-USD price misalignment among the Uniswap (v3) pools on different rollups in Figure 4(a).
+For comparative analysis, we incorporate data from Binance and Uniswap (v3) on Ethereum.
+Figure 4(b) illustrates the MAV that could be achieved within the study period.
+Highest discrepancy.
+The most significant average price discrepancies, approximately 20 basis points (0.0020), are observed between ZKsync and other rollups, which can be attributed to the currently limited token reserves in the Uniswap (v3) liquidity pools on ZKsync, as illustrated in Figure 6 in §0.C.
+These constrained reserves cause the highest price volatility, due to the AMM’s price impact, and restrict arbitrage opportunities.
+Consequently, the arbitrage opportunities between ZKsync and other rollups hovered around $10000$ to $100000$ units during the period we studied.
+Second highest discrepancy.
+The second most significant price differences with Binance are observed on Uniswap on Ethereum.
+This is due to Ethereum’s slower block production rate of about 12 seconds, compared with 0.25 to 2 seconds on rollups.
+The largest arbitrage opportunities are found between Ethereum and Binance, with values reaching up to \$100 million.
+However, cross-rollup arbitrage opportunities exceed \$1 million in Arbitrum, Base, and Optimism and account for about 40 basis points of trading volume, compared to 4 basis points of trading volume on Ethereum.
+Cross-rollup vs.
+DEX-to-CEX.
+The cross-rollup arbitrage opportunities within the analyzed AMM pools on Arbitrum, Base, and Optimism already surpassed the value of arbitrage between these rollups and CEX.
+This can be attributed to the higher price volatility observed in rollup AMMs compared to CEX.
+((a)) Mean (relative) price difference.
+((b)) Maximal Arbitrage Value in USD.
+Figure 4: A comparative analysis of price disparities and arbitrage opportunities across rollups for the WETH-USDC and WETH-USDC.e (e) pools on Uniswap (v3).
+For reference, the Ethereum WETH-USDC Uniswap (v3) pool and Binance are included.
+7 Discussion
+Price disparities between trading venues reveal inefficiencies in financial systems.
+Untapped arbitrage on L2s may result from concerns or misunderstandings of rollup mechanics.
+They can be related to the centralized sequencer, rollup upgradability, bridging risk, or transaction finality.
+DeFi participants, particularly institutional investors, may find arbitrage opportunities less appealing compared to those available on the Ethereum network.
+This is attributed to the lower trading volumes and liquidity reserves observed on rollups.
+Layer-2 MEV.
+Whereas most of the arbitrage transactions on Ethereum are executed using relayer systems such as Flashbots [25, 45], similar mechanisms on Layer-2 (L2) solutions have not yet been developed.
+Furthermore, a significant proportion of MEV on Ethereum is attributed to arbitrage, with approximately 60% being non-atomic arbitrage [12, 16, 24] (between Ethereum DEX and CEX) and 40% being cyclic (atomic) arbitrage within Ethereum.
+Cyclic arbitrage is already present on L2s, but it can be expected that, similar to Ethereum, the majority of arbitrage on L2 will be non-atomic, involving other rollups, Ethereum, or CEX.
+L2 Arbitrage Risk.
+The main risks associated with L2 MEV currently stem from state (price) drift, particularly in cross-rollup arbitrage and CEX-DEX arbitrage.
+This drift means that price changes may occur before both legs of the transaction finalize.
+In contrast to cycling arbitrage, cross-rollup and DEX-CEX arbitrage are not atomic.
+To ensure successful transaction execution, arbitrageurs on L2s employ one of two primary strategies.
+Given that centralized sequencer operators manage most rollups, it is generally assumed that transactions within a block are not reorganized.
+Thus, the predominant strategy for arbitrageurs and MEV searchers on L2s involves partnering with the centralized sequencer to expedite transaction execution.
+Alternatively, another strategy entails saturating the L2 network with a high volume of potential arbitrage transactions.
+Leveraging the lower gas fees on L2s, this approach can be profitable.
+In particular, the rate of reverted transactions in rollups has increased significantly [43], following the Dencun upgrade, which reduced transaction fees.
+Impact of a Shared Sequencer.
+Engaging in cross-rollup arbitrage in the current landscape is fraught with risk.
+By merely submitting the transactions to the designated rollups’ mempools, the arbitrageur is exposed to the risk that, by the time these transactions are sequenced, price fluctuations may occur in one or both rollups.
+Consequently, a subset of these transactions might fail, potentially resulting in financial losses for the arbitrageur.
+The quantification of this risk presents a complex problem that we leave open for future investigation.
+For example, if the sequencing of transactions for both rollups $A$ and $B$ were to be managed by a single entity, the aforementioned risk could be completely mitigated.
+This is because the definitive state of a rollup is inherently determined by the contents of the most recent L1 block.
+Consequently, the proposer of the subsequent L1 block possesses exclusive authority over the progression of a rollup’s state.
+Specifically, the sequencer has the power to exercise discretion over which rollup transactions to include and in what sequence.
+Consider the scenario where following the most recent L1 block, at height $h$, both rollups ($A$ and $B$) exhibit a state that engenders an arbitrage opportunity between their respective Uniswap pools.
+This particular state, and consequently the associated prices, can only be modified through the execution of transactions on the rollups.
+Hence, the sequencer responsible for the L1 block at height $h+1$ solely governs the subsequent progression of the rollup state for block $h+1$.
+Should the same sequencer be tasked with incorporating transactions for both rollups, it can ensure the execution of the arbitrage before any alteration to the rollup state occurs, for instance, by placing the arbitrage transactions at the forefront of the respective rollup’s blocks.
+The principal challenge within the aforementioned process lies in transferring ETH between the two rollups, which typically requires burning of ETH on rollup $A$ and minting on rollup $B$.
+In particular, the procedure of withdrawing from $A$ to L1 and then depositing from L1 to $B$ is likely to be excessively time consuming.
+Consequently, the arbitrageur is compelled to utilize a direct bridge between $A$ and $B$, capable of facilitating the cross-rollup transfer through a minimal number of transactions.
+The bridge, an actor relied upon (or potentially implemented) [^1] The most rudimentary implementation of such a bridge involves maintaining an ETH reserve on the respective rollups.
+by the arbitrageur, holds ETH on both rollup $A$ and rollup $B$.
+This enables the bridge to execute the ETH transfer from $A$ to $B$ for the arbitrageur in just two transactions.
+Single-block finality, ZK bridging, and Flash Loans.
+In the aforementioned scenario, we evaluated a trusted bridge mechanism for transferring ETH from rollup $A$ to rollup $B$.
+This mechanism requires the bridge to 1) maintain adequate ETH reserves on rollup $B$, and 2) authorize the bridging transaction that mints ETH on rollup $B$.
+This stipulation could be obviated if the rollup state were finalized instantaneously, rather than being subject to a delay as is characteristic of all prominent rollups currently in deployment.
+It is critical to clarify that we are addressing the finality of the rollup block relative to L1, rather than the finality of the L1 block itself.
+The assurance sought is that the rollup’s state can only be reverted in the event of an L1 reorganization.
+Now, assume that rollup $A$ operates as a ZK rollup and that a validity proof for a transaction block can be computed in a timely manner, allowing it to be included within the block itself.
+This facilitates the implementation of a bridge via a straightforward smart contract on rollup $B$.
+The arbitrageur can burn the ETH acquired from the swap through a transaction on rollup $A$, and the validity proof of this transaction (now incorporated in the same block) can be directly validated by the bridge contract on rollup $B$, which then mints the equivalent amount of ETH and transfers it to the arbitrageur on rollup $B$.
+We can extend this methodology without additional presuppositions to bridge the USDC (obtained by the arbitrageur from its subsequent swap on rollup $B$) back to rollup $A$, thereby completing the arbitrage cycle and resulting in the arbitrageur accruing more USDC on rollup $A$ than initially invested.
+The atomic execution of the entire arbitrage process presents another opportunity for the arbitrageur: securing the initial investment (USDC) via a flash loan on rollup $A$, implying the addition of a borrowing transaction before the arbitrage and a subsequent corresponding repayment transaction.
+If all transactions necessary for the arbitrage (including the flash loan) are sequenced at the outset of the respective rollups’ blocks, a condition readily guaranteed by a shared sequencer, the arbitrage process is rendered entirely risk-free.
+Ethical MEV.
+There is a debate as to whether rollup sequencers should only allow back-running MEV, as suggested by the time-boost mechanism [34].
+It is crucial to distinguish between front-running and back-running MEVs and their combined effects.
+Front-running MEV, especially with back-running MEV, leads to exploitative attacks like sandwich, imitation, and just-in-time liquidity attacks [48].
+In contrast, back-running MEV mainly supports arbitrage [40] and liquidation [36], which are beneficial for DeFi by promoting price balance (arbitrage) and strengthening lending mechanisms (liquidations).
+Arbitrage measures and LVR.
+The LVR metric compares the profitability of arbitrageurs and LPs, whereas the MAV metric measures the arbitrage profit as net LVR per block.
+Furthermore, LVR assumes the price alignment within one block, so applying it to rollups results in double counting due to multi-block price discrepancies.
+It can be debatable whether arbitrageurs’ profits should be compared with LPs’ rewards.
+LPs earn pro-rata fees based on provided liquidity, while arbitrageurs face competition and uncertainty when executing arbitrage transformations.
+Thus, liquidity provision and arbitrage are usually handled by separate entities due to different infrastructure, software, and expertise.
+As trading activities transition to rollups with Ethereum as a settlement layer, cross-rollup arbitrage is likely to increase, especially when rollup liquidity exceeds that of CEXs and Ethereum’s DEX.
+Arbitrage opportunities, particularly for tokens not listed or liquid on CEXs, may arise between Ethereum and rollups.
+8 Conclusions
+This paper quantifies the potential MEV in rollups that can arise from price disparities among rollups, and between rollups and CEXs.
+We study WETH-USDC, the most traded cryptocurrency pair in the Ethereum ecosystem, and source prices the liquidity pools of Uniswap (v3) on Ethereum and its L2s: Arbitrum, Base, Optimism and ZKsync Era.
+Binance is used as a reference CEX.
+By analyzing price discrepancies after each block, we identify more than 0.5 million unexploited arbitrage opportunities.
+Significantly, these opportunities persist on average for a period of 10 to 20 blocks.
+Thus, to avoid double counting of the same opportunity, we record only the Maximal Arbitrage Value (MAV) in each price misalignment period.
+The estimated MAV on Arbitrum, Base and Optimism pools ranges from 0.03% to 0.05% of trading volume, and oscilates around 0.25% in the newly-deployed Uniswap (v3) WETH-USDC pool on ZKsync Era.
+The empirical MAV approach, filtering repeated price misalignments, results in five-time lower arbitrage profits compared to the Loss-Versus-Rebalancing (LVR) metric.
+We observed that lower-volume swaps are more frequent on rollups compared to Ethereum.
+Swaps occur 2–3 times more often on L2s, but trade volumes are approximately five times lower.
+Block times impact rollup transactions; Ethereum averages one swap per block, while rollups have fewer swaps per block despite higher total swaps.
+Specifically, swaps occur every third block on Base, fifth on Optimism, and tenth on Arbitrum.
+Finally, these insights contribute to ongoing discussions on rollup design and optimization, particularly in the context of MEV auctions that can be held by sequencers.
+Acknowledgements
+The authors would like to thank Matej Pavlovic and Maria Inês Silva for their valuable insights and feedback throughout the course of this research.
+This research article is a work of scholarship and reflects the authors’ own views and opinions.
+It does not necessarily reflect the views or opinions of any other person or organization, including the authors’ employer.
+Readers should not rely on this article for making strategic or commercial decisions, and the authors are not responsible for any losses that may result from such use.
+Appendix 0.A Background on Automated Market Makers
+Figure 5: Architecture of a rollup.
 Decentralized Exchanges (DEXs) facilitate the direct exchange of tokens on the blockchain in a non-custodial and atomic way.
-The implementation of a Limit Order Book (LOD) on a blockchain is impractical due to storage costs and computational complexity.
+The implementation of a Limit Order Book (LOD) on a blockchain is impractical due to storage costs and computational complexity [29].
 Consequently, DEXs employ Automated Market Makers (AMMs) to determine the exchange rate between tokens.
-This exchange rate is derived based on the quantity of tokens held in liquidity pools, which are supplied by Liquidity Providers (LPs).
-Constant Function Market Makers (CFMMs) are the most common type of AMMs.
+This exchange rate is derived based on the quantity of tokens held in liquidity pools, which are supplied by Liquidity Providers (LPs) [29].
+Constant Function Market Makers (CFMMs) are the most common type of AMMs [46].
 They use a reserve curve function $F:\mathbb{R}^{N}_{+}\rightarrow\mathbb{R}$, which maps $N$ token reserves $x_{i}$ to a fixed invariant $L$.
@@
 For simplicity, we refer to them as LP fees.
-Constant Product Market Makers (CPMM).
-Uniswap (v2), one of the first and most popular AMMs, operates as a CPMM [5]. Its trading mechanism is governed by the formula:
-$x_{1}\cdot x_{2}=L,$
+Constant Product Market Makers (CPMM).  Uniswap (v2), one of the first and most popular AMMs, operates as a CPMM [2]. Its trading mechanism is governed by the formula:
+$x_{1}\cdot x_{2}=L,$ (10)
 where $x_{i}$ represent the balances of the two tokens in the liquidity pool, and $L$ is a constant indicating the pool’s liquidity level.
 CPMM with Concentrated Liquidity (CLMM).
-Uniswap v3 introduced concentrated liquidity, an enhancement to CPMMs, to improve the efficiency of the utilization of liquidity within the pool [6].
+Uniswap (v3) introduced concentrated liquidity, an enhancement to CPMMs, to improve the efficiency of the utilization of liquidity within the pool [3].
 This allows LPs to specify a price range to provide liquidity.
 For a two-token pool ($N=2$) within the price range $[p_{j},p_{k}]$, the trade invariant is:
-$\left(x_{1}+\frac{L}{\sqrt{p_{k}}}\right)\left(x_{2}+L\cdot\sqrt{p_{j}}\right)=L^{2},$
+$\left(x_{1}+\frac{L}{\sqrt{p_{k}}}\right)\left(x_{2}+L\cdot\sqrt{p_{j}}\right)=L^{2},$ (11)
 where the variables hold similar meanings as previously described.
-The downside of CLMMs is that LPs must continuously monitor and adjust the price range of their liquidity provision, as only trades within this range earn LP fees.
+The downside of CLMMs is that LPs must continuously monitor and adjust the price range of their liquidity provision, as only trades within this range generate LP fees.
 Thus, LPs require liquidity provision strategies to manage LP positions and effectively minimize gas fees.
 Spot Price.
-The spot exchange rate between tokens can be determined as the slope of the $x_{1}-x_{2}$ curve, illustrated in equations 2.2 and 2.2), with partial derivatives of the curve function $F$.
-2.3 Cost of Swapping
+The spot exchange rate between the tokens can be determined as the slope of the $x_{1}-x_{2}$ curve, illustrated in equations 10 and 11, with partial derivatives of the curve function $F$.
+0.A.1 Cost of Swapping
 The total cost of executing swap transactions on an AMM includes gas costs and explicit and implicit DEX fees [46, 23]. For clarity, we express this as follows.
@@
 It is directly related to the size of the trade and the amount of liquidity available in the AMM’s pool.
-3 Data Collection
-We analyze swap data from Ethereum and its roll-ups with the highest trading volumes.
-This includes Arbitrum, Base, and Optimism (optimistic rollups), as well as zkSync Era (ZK-rollup).
-For each blockchain, we focus on the most actively traded pool, WETH-USDC on Uniswap (v3), during the period from December 31st, 2023, to April 30th, 2024.
-Although the Uniswap (v3) deployment on zkSync Era does not have a graphical user interface, users can interact with it via DEX aggregators such as Oku Analytics.
-The data set is sourced from blockchain archive nodes provided by Nansen.
-The number of analyzed swaps, transactions, and block ranges is detailed in Table 1.
-Using the event logs of the Swap method, we recalculate the historical spot prices and liquidity in ticks within the USDC-WETH pools on Uniswap (v3).
-For Uniswap (v3), the spot price after the swap is obtained from sqrtPriceX96 in the event logs.
-Market data for the ETH-USDC exchange rate on CEX are obtained from Binance.
-We examine 1-second Binance closing prices, noting that block production on Ethereum occurs every 12 seconds, while it is around 1–2 seconds on Base, Optimism, and zkSync Era.
-On Arbitrum, block production can be as fast as 0.25 seconds.
-Thus, for each blockchain, except for Arbitrum, computations are performed at the block level.
-For Arbitrum, we analyze the last block in a second.
-The spot price on DEX is determined by the last swap within the block.
-Chain Swaps Transactions Blocks Block Range Ethereum $761005$ $749818$ $475409$ 1890889618908896189088961890889619771559197715591977155919771559 Arbitrum $2400000$ $2367361$ $1709619$ 187373628187373628187373628187373628206540031206540031206540031206540031 Arbitrum (USDC.e) $2258469$ $2232543$ $1648839$ 165788868165788868165788868165788868206540037206540037206540037206540037 Base $1687530$ $1636196$ $1145789$ 863892986389298638929863892913866123138661231386612313866123 Optimism (USDC.e) $1186780$ $1136839$ $902789$ 114234215114234215114234215114234215119461410119461410119461410119461410 zkSync Era (USDC.e) $46417$ $46379$ $45364$ 2290992322909923229099232290992332843035328430353284303532843035
-Table 1: List of analyzed WETH-USDC and WETH-USDC.e liquidity pools on Uniswap (v3) on Ethereum and its rollups, spanning the period from 31st December 2023 to 30th April 2024.
-Native USDC and Bridged USDC.e.
-USDC is a fiat-backed stablecoin issued by Circle.
-It can exist on rollups either as a native token or a bridged one.
-The native USDC is directly supported by Circle’s off-chain reserves, while USDC.e is a token bridged from Ethereum.
-Generally, USDC.e is introduced to the blockchain first, followed by the issuance of the native token.
-In most rollups, USDC and USDC.e coexist for legacy purposes.
-The native USDC is resistant to hacking and can be easily transferred across blockchains using the Circle bridge.
-4 Swap Dynamics
-Tables 2 and 3 present the characteristics of swaps on Ethereum and its rollups.
-As shown in Table 2, swap transactions on rollups have a lower volume compared to those on Ethereum, although they occur more frequently.
-On average, there are 2–3 times more swaps on rollups, but the trading volume is about five times lower.
-Consequently, the median swap volume on Ethereum is $4239.45$ USD, while it is $2201.23$ USD on Arbitrum, $173.46$ USD on Base, $331.28$ USD on Optimism, and only $20$ USD on zkSync Era.
-Table 3 explains the effect of block times on rollup swap transactions.
-Despite having a higher number of swaps, there are fewer swaps per block on rollups because of faster block production.
-In Ethereum, there are, on average, $0.88$ swaps in every block, with every block containing swap transactions having 1.6 swap events.
-In contrast, every third block on Base, fifth block on Optimism, and tenth block on Arbitrum contain a swap.
-Furthermore, there are fewer swap events per block with swap transactions (approximately 1–1.47) on rollups.
-On average, there are slightly more swap events per transaction (1.01) on both rollups and Ethereum, indicating that some swaps are performed in batches.
-This could imply the exploitation of MEV by certain wallets or decentralized exchange (DEX) aggregators initiating swaps on behalf of users.
-Chain Swaps Total Volume mn Avg Swap Volume Std Swap Volume Median Swap Volume
-Ethereum $761005$ $35175$ $46222.26$ $142890.62$ $4239.45$
-Arbitrum $2400000$ $7266$ $3027.66$ $6738.06$ $1125.01$
-Arbitrum (USDC.e) $2258469$ $8159$ $3612.86$ $5976.01$ $2201.23$
-Base $1687530$ $2996$ $1775.55$ $5611.25$ $173.46$
-Optimism (USDc.e) $1186780$ $1457$ $1228.30$ $2842.73$ $331.28$
-zkSync (USDC.e) $46417$ $8.01$ $172.67$ $426.52$ $20.99$
-Table 2: An analysis of the absolute swap volumes within the WETH-USDC and WETH-USDC.e liquidity pools on Uniswap (v3) across the Ethereum network and its rollups, spanning the period from 31st December 2023 to 30th April 2024.
-Chain Block Time (s) Swaps per Block Swaps per Transaction Swaps Per Block with Transaction Ethereum 12.12 $0.88$ $1.01$ $1.60$ Arbitrum 0.25 $0.13$ $1.01$ $1.40$ Arbitrum (USDC.e) 0.25 $0.06$ $1.01$ $1.37$ Base 2.00 $0.32$ $1.03$ $1.47$ Optimism (USDC.e) 2.00 $0.23$ $1.04$ $1.31$ zkSync Era (USDC.e) 1.05 $0.00$ $1.00$ $1.02$
-Table 3: An analysis of the swaps per transactions and blocks within the WETH-USDC and WETH-USDC.e liquidity pools on Uniswap (v3) on the Ethereum network and its rollups, spanning the period from 31st December 2023 to 30th April 2024.
-5 Theory of Arbitrage and LVR
-Price disparities across different trading platforms (both DEXs and CEXs) present arbitrage opportunities.
-In this section, we calculate the transaction volume required to align prices between AMM-DEX and CEX, given the existing token reserves in an AMM liquidity pool.
-Trading this volume maximizes the arbitrageur’s profit, which is called net LVR [24], or the Maximal Arbitrage Value (MAV) [13].
-For simplicity, we further assume that trading on the CEX has no price impact due to the typically higher reserves compared to AMM-DEXs.
-Maximal Arbitrage Value.
-Consider a scenario in which there is a price discrepancy between a CEX and an AMM, the CEX having a substantially higher market depth, resulting in zero market impact.
-Let $P_{c}$ and $P_{a}$ be the prices at CEX and AMM, respectively, and assume that $P_{c}$ < $P_{a}$.
-Let $V_{max}$ denote the maximum trade volume that was bought on the CEX for $P_{c}$ and sold on the AMM for $P_{a}$ equals the prices.
-With $V_{max}$ being the trading volume that equals the prices at AMM and CEX and percentage price impact $\rho(V_{max})$, MAV is defined as:
-$MAV=V_{max}\cdot\Big{(}P_{a}\cdot(1-\rho(V_{max}))-P_{c}\Big{)},$
-or equivalently
-$MAV=V_{max}\cdot(P_{a}-P_{c})-V_{max}\cdot P_{a}\cdot\rho(V_{max}).$ (1)
-Arbitrage between a CPMM and CEX.
-Let us assume that we have two tokens $X$ and $Y$, the respective reserves in the liquidity pool of AMM are $x_{1}=x$ and $x_{2}=y$.
-Assuming that we are dealing with a CPMM, the act of trading follows the $x_{1}\cdot x_{2}=x\cdot y=L$ formula, and the instantaneous exchange rate $P_{a}$ for an infinitesimal trade is
-$P_{a}=\frac{x}{y}.$ (2)
-and the percentage price impact on a CPMM (such as Uniswap v2) is given by $\Delta x/x$ for a swap of $\Delta x$ for $\Delta y$. the MAV can be expressed as:
-$MAV=V_{max}\cdot(P_{a}-P_{c})-\frac{V_{max}^{2}P_{a}}{x},$ (3)
-Following [16, 13] we can compute the first derivative of Eq. (3) with respect to $V_{max}$ to find the max points, receiving:
-$V_{max}=\frac{y\cdot(P_{a}-P_{c})}{2P_{a}}.$ (4)
-And, by substituting (4) into (3) the MAV is written as:
-$MAV=\frac{y\cdot(P_{a}-P_{c})^{2}}{4P_{a}}.$ (5)
 Arbitrage between a CLMM and CEX.
@@
 $MAV_{Uniswap\_v3}=\sum_{i:P_{a}\in[p(i),p(i+1)]}^{i:P_{c}\in[p(i),p(i+1)]}MAV_{i}.$ (16)
-6 Empirical Results
-When price discrepancies arise between trading platforms, arbitrageurs can take advantage of these opportunities to profit from price corrections.
-This activity benefits both the arbitrageurs, who gain from the price differences, and traders, who enjoy promptly adjusted and accurate exchange rates.
-The only parties that might suffer losses from price corrections are LPs at AMM-DEX.
-As a result, arbitrage profits are compared with LP losses in the loss-versus-rebalancing (LVR) metric [24].
-In this section, we examine the price differences between WETH-USDC (and USDC.e) AMM pools on Ethereum and its rollups — Arbitrum, Base, Optimism, zkSync Era — and Binance, the leading cryptocurrency CEX.
-We measure the decay time of these discrepancies and quantify the resulting arbitrage opportunities using the proposed Maximal Arbitrage Value (MAV) metric, similar to the net LVR.
-We begin by presenting a methodology to estimate the empirical MAV, which accounts for the multi-block decay time of price disparities and avoids the double-counting of the same opportunities.
-6.1 Methodology
-To prevent double-counting identical arbitrage opportunities that span several blocks, we adopt the methodology outlined in the empirical MAV assessments [13] and modify it for block-by-block analysis.
-It includes the following components:
-  1. 1.
-    Block Intervals: We analyze Binance ETH-USDC closing prices every second, and compare them with the spot price on the AMM-DEX at every block.
-The spot price on the DEX is determined based on the last swap within the block.
-The block production on Ethereum takes approximately 12 seconds, followed by around 2 seconds on Base, Optimism, and zkSync Era.
-On Arbitrum, blocks can be produced more quickly, approximately every 0.25 seconds.
-In such cases, we analyze the last block within the second.
-Figure 3(a) shows the ETH-USDC exchange rates on Binance and zkSync Era, highlighting the existence of price differences.
-  2. 2.
-    Price Re-Alignment Threshold: We only consider the price deviations that surpass a set threshold.
-This threshold is determined from the empirical distribution of price differences (see Fig.
-3(b)), mainly focusing on outlier values (i.e., those beyond 1.5 times the interquartile range above the third quartile).
-This selected threshold ensures that prices are generally aligned but allows for adjustments to enhance flexibility.
-  3. 3.
-    Empirical MAV: Figure 3(c) illustrates the extent of price discrepancies.
-If a price difference continues for a prolonged duration of multiple blocks, we document only the highest MAV within each time segment of price discrepancy until re-alignment happens.
-This approach avoids counting the same ongoing price discrepancy multiple times.
-  4. 4.
-    Decay Time: Once all instances of empirical MAV (i.e., the peak MAV during periods of price misalignment) are identified, we determine the duration required for prices to re-align (i.e., when the magnitude of the price difference drops below the specified threshold).
-((a)) Time series of prices for USDC-ETH.
-((b)) Distribution of deltas.
-((c)) Price differences at the block level, with the red dashed line indicating the threshold at which prices are assumed to have re-aligned.
-Missing bars represent blocks where no trading activity occurred on the AMM.
-Figure 3: These plots illustrate the price differences for USDC-ETH between the AMM on zkSync and Binance.
-The AMM spot price fluctuates around the CEX price, showing how price misalignments first widen and subsequently narrow over a period of several minutes.
-6.2 Empirical MAV from WETH-USDC pool
-As stated previously, greater chances for arbitrage profit arise when there is a significant price disparity.
-Nevertheless, the possible profit is also dependent on the overall trade volume and the pool’s reserves.
-Chain MAV % Volume Net LVR % Volume
-Ethereum $52488479.88$ $0.15$ $126531170.12$ $0.36$
-Arbitrum $2699032.60$ $0.04$ $20686037.49$ $0.28$
-Arbitrum (USDC.e) $2707008.96$ $0.03$ $19332648.22$ $0.24$
-Base $1293771.96$ $0.04$ $32663704.76$ $1.09$
-Optimism (USDC.e) $690868.44$ $0.05$ $3893401.00$ $0.27$
-zkSync Era (USDC.e) $20110.31$ $0.25$ $96016.54$ $1.20$
-Table 4: Assessment of arbitrage opportunities within the WETH-USDC and WETH-USDC.e liquidity pools on Uniswap (v3) on Ethereum and its rollups from 31 December 2023 to 30 April 2024.
-The metrics encompass the total Maximal Arbitrage Value (MAV) and net Loss-Versus-Rebalancing (LVR), both in absolute terms and as percentages of trading volume.
-Table 4 shows the total arbitrage opportunities (MAV) during the study periods, expressed as an absolute number and as a percentage of the trading volume.
-The greatest price differences occur on Ethereum, resulting in these arbitrage opportunities.
-However, these opportunities cannot be fully realized due to gas fees and MEV auction fees, which account for about 0.15% of the trading volume.
-The detected arbitrage opportunities on the Ethereum pool may remain untapped due to specific details in the Binance order book, which we cannot assess based on the 1-second data intervals.
-The behavior seen in the rollups shows significant variations from those on Ethereum.
-In particular, MAV in the Arbitrum, Base, and Optimism pools ranges from 0.03% to 0.05% of trading volume, while in the zkSync pool, it is around 0.25%, suggesting that price discrepancies are not yet fully addressed in that rollup.
-Additionally, as expected, empirical MAV, which filters out identical arbitrage opportunities across multiple blocks, leads to lower rewards compared to cumulative net LVR.
-The LVR ranges from 0.24% to 1.2%, which is roughly five times higher than MAV.
-Chain Points Avg MAV $V_{max}$ Avg Decay (s)
-Ethereum $71135$ $737.87$ $499667.10$ 31
-Arbitrum $137162$ $19.68$ $56476.35$ 6.9
-Arbitrum (USDC.e) $164080$ $16.50$ $41370.11$ 8.8
-Base $75729$ $17.08$ $22620.17$ 420
-Optimism (USDC.e) $113428$ $6.09$ $8477.60$ 19
-zkSync Era (USDC.e) $3970$ $5.07$ $1966.76$ 370
-Table 5: Assessment of arbitrage opportunities within the WETH-USDC and WETH-USDC.e liquidity pools on Uniswap (v3) on Ethereum and its rollups from December 31st, 2023 to April 30th, 2024.
-The metrics encompass the quantity of identified arbitrage opportunities, mean Maximal Arbitrage Value (MAV), average transaction volume yielding maximal arbitrage, and mean decay time of the price disparities.
-Table 5 provides further insight into the opportunities analyzed.
-The highest quantity of arbitrage opportunities is observed in the two Arbitrum pools, which can be attributed to higher block production rates.
-Base and Optimism also exhibit more arbitrage opportunities compared to the Ethereum pool.
-In particular, the average MAV is significantly lower on rollups, approximately 5–20 USD compared to over 700 USD on Ethereum.
-However, the transaction volume required to achieve this arbitrage is also lower (approximately 2.0002.0002.0002.0005.0005.0005.0005.000 USD on rollups and 0.5 million USD on Ethereum).
-The average decay on rollups varies: for Arbitrum, it is approximately 7–9 seconds (with a 0.25-second block time), followed by 19 seconds on Optimism (2-second block time), 370 seconds on zkSync (1.05-second block time), and 420 seconds on Base (2-second block time).
-In each instance, the spans cover 10–20 blocks or more.
-Figure 4: Empirical daily analysis of Maximal Arbitrage Value (MAV), price discrepancies, and liquidity in the current tick within the WETH-USDC liquidity pool on Uniswap (v3) deployed on the Arbitrum network.
-Figure 4 illustrates the changes in price over time, the observed Maximal Arbitrage Value (MAV), and the pool’s virtual reserves within the current tick for the most active pool analyzed, WETH-USDC in Arbitrum.
-It specifically shows the maximum daily price difference, the daily total of observed MAV, and the end-of-day liquidity in the current pool.
-The data reveal that the daily MAV ranges from \$$2000$ to \$$140000$, with the lowest point occurring during low trading volumes between March and April 2024.
-The smaller price differences with Binance during this period result in a lower MAV.
-Importantly, the days with the highest daily MAV align with those with the largest price differences.
-Figure 5 in the Appendix provides an analysis of liquidity pools on Ethereum, Base, Arbitrum, and Optimism.
-In particular, even with the introduction of native USDC on Arbitrum and its corresponding pool on Uniswap (v3), trading activity and associated arbitrage opportunities remain stable for both native USDC and bridged USDC tokens.
-A significant correlation between arbitrage opportunities and the size of reserves and liquidity within the current tick is observed in the zkSync Era pool.
-The highest arbitrage opportunities occurred on days when liquidity within the current tick was at its maximum.
-In contrast, on Optimism, arbitrage opportunities are closely linked to price differences.
-This can be explained by the relatively small size of the pool on zkSync Era.
-Despite the rise in reserves on Base, there is no corresponding increase in arbitrage profits.
-This indicates that the estimation of arbitrage opportunities is complex and should be evaluated based on price discrepancies, fees, and pool liquidity.
-7 Discussion
-Arbitrage Presence.
-Arbitrage reveals financial system inefficiencies.
-For AMMs on rollups, these may arise from security, infrastructure issues, or lack of understanding of rollup mechanics.
-Untapped arbitrage in rollups can be due to the absence of MEV auctions, low confidence in rollups’ upgradeability, and concerns about transaction finality.
-Misunderstandings about soft and hard finality at roll-up often lead to incorrect assumptions about bridging token durations.
-DeFi asset managers that focus on MEV auctions on Ethereum may avoid rollups due to limited trading volumes and liquidity, restricting arbitrage opportunities on roll-ups.
-LVR as an Arbitrage Measure.
-The main goal of the LVR metric is to compare the profitability between arbitrageurs and LPs rather than to measure arbitrage itself.
-In this research, we evaluate the absolute value of arbitrage opportunities, also known as net LVR (per block).
-As mentioned earlier, LVR assumes the immediate exploitation of arbitrage opportunities, and its direct application to rollups, where price discrepancies last across multiple blocks, would lead to double counting.
-There is an ongoing debate about whether arbitrageurs should be compared to LPs.
-LPs earn their rewards from fees on a pro-rata basis according to the liquidity they provide to the pool.
-In contrast, arbitrageurs are not assured of receiving arbitrage rewards, as they compete with each other to exploit the same opportunities.
-Therefore, liquidity provision and arbitrage are typically carried out by different asset management firms.
-Given the distinct infrastructure, software, and expertise needed, it is unlikely that a single asset management firm engages in both liquidity provision and arbitrage extraction.
-MEV on Rollups.
-As roll-up sequencers become decentralized, it can be anticipated that MEV auctions will be introduced to rollups.
-Currently, assuming trust in the centralized sequencer operators to prevent transaction reordering, MEV exploitation is limited to DEXs through back-running mechanisms within the smart contract.
-The back-running MEV can generate additional revenue for the DEX and, if redistributed, can also benefit the LPs contributing to the AMM liquidity pool.
-However, as previously mentioned, this arbitrage exploitation could be carried out by DEXs or specialized entities rather than LPs.
-Positive and Negative MEV.
-The ongoing discussion about the role of MEV, often referred to as MEV attacks due to their negative implications, continues.
-It is crucial to distinguish between front- and back-running MEVs, as well as their combined effects.
-Front-running MEV, especially when paired with back-running MEV, leads to various exploitative strategies, such as sandwich, imitation, and just-in-time liquidity attacks [32].
-However, back-running MEV mainly supports arbitrage extraction and liquidation processes.
-Both arbitrage and liquidation play beneficial roles within DeFi ecosystems, fostering price balance and fair trading (arbitrage) and strengthening lending mechanisms (liquidations).
-As a result, there is a debate on whether rollup sequencers should allow only back-running MEV, as proposed by the time-boost mechanism [22].
-Cross-Rollup Arbitrage.
-Further research should explore cross-rollup arbitrage, especially considering the proposed use of shared sequencers across multiple roll-ups, which would enable atomic transactions between rollups [23].
-Currently, arbitrage occurs mainly between AMMs and CEXs such as Binance, where trading volumes are highest, raising questions about the tokens to be exchanged across rollups.
-Arbitrage opportunities might emerge between Ethereum and rollups, particularly for tokens not listed on CEXs or with lower liquidity on CEXs.
-As trading shifts from Ethereum to roll-ups, with Ethereum increasingly acting as a settlement layer, cross-rollup arbitrage could become more common, especially when rollup liquidity surpasses that of CEXs and Ethereum’s AMMs.
-8 Conclusions
-This paper examines the DeFi dynamics in the Layer-2 ecosystem, focusing on AMMs and arbitrage on Ethereum rollups.
-We analyze swaps and arbitrage on the WETH-USDC pools on Uniswap (v3) on Ethereum, Arbitrum, Base, Optimism, and zkSync Era, noting significant user activity shifts from Ethereum to these rollups.
-By analyzing the price discrepancies between Uniswap and Binance, we identified more than 0.5 million unexploited arbitrage opportunities on rollups.
-Significantly, these opportunities persisted on average for a period of 10 to 20 blocks.
-Thus, we revised the LVR metric to avoid double-counting arbitrage, recording only the highest arbitrage value (MAV) during sustained price differences over multiple blocks.
-MAV in Arbitrum, Base, and Optimism pools ranges from 0.03% to 0.05% of trading volume, while zkSync Era is around 0.25%, indicating unresolved price discrepancies.
-The empirical MAV approach, filtering repeated arbitrage, suggests that LVR metrics overestimated arbitrage by a factor of five.
-We found that lower-volume swaps are more frequent on roll-ups compared to Ethereum.
-Swaps occur 2–3 times more often on rollups, but trade volumes are about five times lower.
-Block times impact rollup transactions; Ethereum averages one swap per block, while rollups have fewer swaps per block despite higher total swaps.
-Specifically, swaps occur every third block on Base, fifth on Optimism, and tenth on Arbitrum.
-In conclusion, our research highlights unique behaviors and opportunities in the Layer-2 AMM-DEX ecosystem.
-Rollups offer not only lower gas fees, but also faster block production, leading to significant differences in the trading and arbitrage dynamics from that of Ethereum.
-These insights contribute to ongoing discussions on rollup design and optimization, particularly in the context of MEV auctions at sequencers.
-Acknowledgements
-This research article is a work of scholarship and reflects the authors’ own views and opinions.
-It does not necessarily reflect the views or opinions of any other person or organization, including the authors’ employer.
-Readers should not rely on this article for making strategic or commercial decisions, and the authors are not responsible for any losses that may result from such use.
-One of the authors’ research was supported by the EPSRC CDT in Mathematics of Random Systems (EPSRC Grant EP/S023925/1).
-Appendix 0.A Appendix
 ((a)) Ethereum
@@
 ((f)) ZKsync Era (USDC.e)
-Figure 5: Empirical Maximal Extractable Value (MAV), price misalignment with Binance, and liquidity within the spot price tick of the WETH-USDC pool on Uniswap (v3) across Ethereum and its rollups from December 31st, 2023, to April 30th, 2024.
+Figure 6: Empirical Maximal Extractable Value (MAV), price misalignment with Binance, and liquidity within the spot price tick of the WETH-USDC pool on Uniswap (v3) across Ethereum and its rollups from December 31st, 2023 to April 30th, 2024.
+Appendix 0.B Dynamics and Cost of Swapping on L2s
+Chain Swaps Total Volume mn Avg Swap Volume Std Swap Volume Median Swap Volume
+Ethereum $761005$ $35175$ $46222.26$ $142890.62$ $4239.45$
+Arbitrum $2400000$ $7266$ $3027.66$ $6738.06$ $1125.01$
+Arbitrum (USDC.e) $2258469$ $8159$ $3612.86$ $5976.01$ $2201.23$
+Base $1687530$ $2996$ $1775.55$ $5611.25$ $173.46$
+Optimism (USDc.e) $1186780$ $1457$ $1228.30$ $2842.73$ $331.28$
+ZKsync (USDC.e) $46417$ $8.01$ $172.67$ $426.52$ $20.99$
+Table 4: An analysis of the absolute swap volumes within the WETH-USDC and WETH-USDC.e liquidity pools on Uniswap (v3) across the Ethereum network and its rollups, spanning the period from 31st of December 2023 to 30th of April 2024.
+Tables 4 and 5 present the characteristics of WETH-USDC swaps on rollups, while Table  6 breaks down the costs of swapping.
+The corresponding values of Ethereum are provided as a benchmark.
+As shown in Table 4, swap transactions on rollups have a lower volume compared to those on Ethereum, although they occur more frequently.
+On average, there are 2–3 times more swaps on rollups, but the trading volume is about five times lower.
+Consequently, the median swap volume on Ethereum is $4239.45$ USD, while it is $2201.23$ USD on Arbitrum, $173.46$ USD on Base, $331.28$ USD on Optimism, and only $20$ USD on ZKsync Era.
+Table 5 explains the effect of block times on rollup swap transactions.
+Despite having a higher number of swaps, there are fewer swaps per block on rollups because of faster block production.
+In Ethereum, there are, on average, $0.88$ swaps in every block, with every block containing swap transactions having 1.6 swap events.
+In contrast, every third block on Base, fifth block on Optimism, and tenth block on Arbitrum contain a swap.
+Furthermore, there are fewer swap events per block with swap transactions (approximately 1–1.47) on rollups.
+On average, there are slightly more swap events per transaction (1.01) on both rollups and Ethereum, indicating that some swaps are performed in batches.
+This could imply the exploitation of MEV by certain wallets or decentralized exchange (DEX) aggregators initiating swaps on behalf of users.
+Table 6 shows the costs of swapping on L2s before and after the Dencun upgrade.
+Although the otal costs and average gas fees decreased, the percentage costs of swaps often remained unchanged or increased.
+This is due to the smaller swap sizes on rollups post-upgrade.
+Chain Block Time (s) Swaps per Block Swaps per Transaction Swaps Per Block with Transaction Ethereum 12.12 $0.88$ $1.01$ $1.60$ Arbitrum 0.25 $0.13$ $1.01$ $1.40$ Arbitrum (USDC.e) 0.25 $0.06$ $1.01$ $1.37$ Base 2.00 $0.32$ $1.03$ $1.47$ Optimism (USDC.e) 2.00 $0.23$ $1.04$ $1.31$ ZKsync Era (USDC.e) 1.05 $0.00$ $1.00$ $1.02$
+Table 5: An analysis of the swaps per transactions and blocks within the WETH-USDC and WETH-USDC.e liquidity pools on Uniswap (v3) on the Ethereum network and its rollups, spanning the period from the 31st of December 2023 to the 30th of April 2024.
+Chain Total Fee Gas Fee LP Fee Slippage and Price Impact
+Whole Period:
+Ethereum $97.20$ (21bps) $37.13$ (8bps) $23.11$ (5bps) $65.33$ (14bps)
+Arbitrum $2.46$ (8bps) $0.65$ (2bps) $1.51$ (5bps) $1.82$ (6bps)
+Arbitrum (USDC.e) $5.55$ (15bps) $3.41$ (9bps) $1.81$ (5bps) $2.18$ (6bps)
+Base $1.76$ (10bps) $0.52$ (3bps) $0.89$ (5bps) $1.24$ (7bps)
+Optimism (USDC.e) $1.33$ (11bps) $0.31$ (2bps) $0.61$ (5bps) $1.10$ (9bps)
+ZKsync (USDC.e) $0.87$ (50bps) $0.25$ (15bps) $0.52$ (30bps) $0.65$ (38bps)
+Pre-Blobs:
+Ethereum $92.96$ (22bps) $39.30$ (9bps) $20.94$ (5bps) $59.43$ (14bps)
+Arbitrum $9.68$ (10bps) $4.16$ (4bps) $4.73$ (5bps) $5.76$ (6bps)
+Arbitrum (USDC.e) $6.28$ (11bps) $3.08$ (6bps) $2.77$ (5bps) $3.30$ (6bps)
+Base $0.43$ (22bps) $0.11$ (6bps) $0.10$ (5bps) $0.36$ (18bps)
+Optimism (USDC.e) $1.55$ (10bps) $0.36$ (2bps) $0.78$ (5bps) $1.27$ (8bps)
+ZKsync (USDC.e) $1.37$ (49bps) $0.37$ (13bps) $0.83$ (30bps) $1.03$ (37bps)
+Post-Blobs:
+Ethereum $102.26$ (20bps) $34.56$ (7bps) $25.69$ (5bps) $72.35$ (14bps)
+Arbitrum $2.00$ (8bps) $0.42$ (2bps) $1.31$ (5bps) $1.57$ (6bps)
+Arbitrum (USDC.e) $5.05$ (22bps) $3.64$ (16bps) $1.15$ (5bps) $1.41$ (6bps)
+Base $1.99$ (10bps) $0.59$ (3bps) $1.02$ (5bps) $1.40$ (7bps)
+Optimism (USDC.e) $1.17$ (12bps) $0.26$ (3bps) $0.49$ (5bps) $0.97$ (10bps)
+ZKsync (USDC.e) $0.32$ (56bps) $0.12$ (21bps) $0.17$ (30bps) $0.24$ (40bps)
+Table 6: Breakdown of swap fees at the WETH-USDC and WETH-USDC.e pools at Uniswap (v3) at Ethereum and its rollups during the period 12.31-30.04, 13.03-30.04.24 (pre-blobs) and 14.3-30.04.24 (post-blobs).
+Appendix 0.C Empirical Maximal Extractable Value (MAV)
+Figure 6 shows an analysis of MAV across Ethereum, Base, Arbitrum, and Optimism.
+Notably, despite the introduction of native USDC on Arbitrum and its corresponding pool on Uniswap (v3), trading activity and arbitrage opportunities have remained steady for both native and bridged USDC tokens.
+In the ZKsync Era pool, a strong relationship is observed between arbitrage opportunities and the size of reserves and liquidity within the current tick.
+The most significant arbitrage opportunities occurred when liquidity within the current tick reached its peak.
+Conversely, on Optimism, arbitrage opportunities are more closely tied to price discrepancies, likely due to the relatively small pool size on ZKsync Era.
+Interestingly, while reserves have increased on Base, this hasn’t led to a proportional rise in arbitrage profits.
+This suggests that arbitrage potential is complex and must be evaluated not only by price differences but also by considering fees and pool liquidity.